### v1

# From Zipf‘s Law to Neural Scaling through Heaps‘ Law and Hilberg‘s Hypothesis

From Zipf‘s Law to Neural Scaling through
Heaps‘ Law and Hilberg‘s Hypothesis
Łukasz Dębowski Łukasz Dębowski is with the Institute of Computer Science, Polish Academy of Sciences, ul. Jana Kazimierza 5, 01-248 Warszawa, Poland, e-mail: ldebowsk@ipipan.waw.pl.
We inspect the deductive connection between the neural scaling law and Zipf‘s law—two statements discussed in machine learning and quantitative linguistics. The neural scaling law describes how the cross entropy rate of a foundation model—such as a large language model—changes with respect to the amount of training tokens, parameters, and compute. By contrast, Zipf‘s law posits that the distribution of tokens exhibits a power law tail. Whereas similar claims have been made in more specific settings, we show that the neural scaling law is a consequence of Zipf‘s law under certain broad assumptions that we reveal systematically. The derivation steps are as follows: We derive Heaps‘ law on the vocabulary growth from Zipf‘s law, Hilberg‘s hypothesis on the entropy scaling from Heaps‘ law, and the neural scaling from Hilberg‘s hypothesis. We illustrate these inference steps by a toy example of the Santa Fe process that satisfies all the four statistical laws.
Key words: neural scaling law, Zipf‘s law, Heaps‘ law, Hilberg‘s hypothesis, Santa Fe processes
MSC 2020: 94A17, 60G10
1 Introduction
It has been increasingly recognized in the machine learning literature [50, 61, 64, 14, 70, 71] that the neural scaling law observed for contemporary foundation models—such as large language models—may arise as a consequence of Zipf‘s law or similar distributional regularities of natural language. This emerging perspective suggests that some remarkable empirical regularities in large-scale deep learning need not be explained solely by architectural, optimization, or hardware considerations, but instead reflect intrinsic statistical constraints of natural texts—human linguistic data.
In the present paper, we aim to provide an actually simple but systematic derivation of this chain of implications that extends earlier results in information theory and quantitative linguistics [23, 25, 26, 28]. We will prove formal theorems about as general stochastic processes as possible rather than experiment with particular empirical data. Our goal is to consolidate knowledge of several scientific disciplines by means of mathematical deduction.
The conceptual trajectory is as follows. Beginning with Zipf‘s law for word frequencies, we derive Heaps‘ law for vocabulary growth. From Heaps‘ law we extract a formal version of Hilberg‘s hypothesis on the sublinear growth of block entropy. Finally, we show that Hilberg‘s hypothesis leads to the neural scaling that ties model performance to training data size, model size, and amount of training compute. In parallel, we will discuss an example of Santa Fe processes—toy stochastic sources that exhibit these statistical laws. Each link in this chain introduces its own assumptions and modeling choices. By isolating these steps, we hope to illuminate where the derivations rely on idealizations and how they might be strengthened.
Notations.
Notation $(X_{t})_{t\in\mathbb{Z}}$ denotes an integer-time countable-alphabet stochastic process, $X_{j}^{k}:=(X_{j},X_{j+1},\ldots,X_{k})$ is a string of random tokens, and $\mathcal{X}_{j}^{k}:=\left\{X_{j},X_{j+1},\ldots,X_{k}\right\}$ is a random set of types. We adopt that $X_{1}^{0}$ is the empty string and $\mathcal{X}_{1}^{0}$ is the empty set. Notation $H(X):=\operatorname{\mathbf{E}}(-\log P(X))$ stands for the Shannon entropy of a random variable and $V(\mathcal{X}):=\operatorname{\mathbf{E}}\operatorname{\#}\mathcal{X}$ denotes the expected cardinality of a random set. The conditional entropy is H(X|Q):=H(X,Q)H(Q)H(X|Q):=H(X,Q)-H(Q), the mutual information is I(X;Y):=H(X)H(X|Y)I(X;Y):=H(X)-H(X|Y), and the conditional mutual information is I(X;Y|Q):=H(X|Q)H(X|Y,Q)I(X;Y|Q):=H(X|Q)-H(X|Y,Q). We also use notations
$f(x)\stackrel{{\scriptstyle*}}{{<}}g(x)$ $\iff f(x)\leq cg(x)\text{ for a $c>0$},$ (1)
$f(x)\stackrel{{\scriptstyle*}}{{>}}g(x)$ $\iff f(x)\geq cg(x)\text{ for a $c>0$},$ (2)
$f(x)\stackrel{{\scriptstyle*}}{{=}}g(x)$ $\iff f(x)\stackrel{{\scriptstyle*}}{{<}}g(x)\land f(x)\stackrel{{\scriptstyle*}}{{>}}g(x).$ (3)
Statistical laws at large.
Let us briefly present the five concepts to be related in our reasoning—in the chronological order of their discovery.
Zipf‘s law. The oldest known of quantitative linguistic laws, Zipf‘s law asserts that, for texts in natural language, the frequency of the $k$-th most common type of word decays approximately as $k^{-\alpha}$, where $\alpha\approx 1$. This regularity was noticed over one hundred years ago [31, 17, 90]. An empirical study of this law across one hundred languages can be found in [63]. Variants of Zipf‘s law with double regimes are also well known [34, 69, 72, 39, 86]. Similar power-law distributions are observed also in ecology, sociology, economics, and physics [91], being a hallmark of complex systems. The literature on Zipf‘s law is scattered over diverse venues. Historical references can be found in [59]. The departure point of a statistically informed theory is [52, 4].
In spite of sheer literature coverage, there is no single explanation of Zipf‘s law. The law can be explained by diverse mechanisms ranging from monkey-typing [62, 68], through preferential attachment [80] and game theory [41], to potential links with semantics and information theory [26]. What matters for our purposes is that stochastic processes $(K_{t})_{t\in\mathbb{Z}}$ over natural numbers with the Zipf probability distribution
$P(K_{t}=k)\stackrel{{\scriptstyle*}}{{=}}k^{-\alpha},\quad\alpha>1,$ (4)
provide a compact description of heavy-tailed data distributions. Observation (4) will be our departure point.
Heaps‘ law. A closely related law, Heaps‘ law, also known as Herdan‘s law, describes vocabulary growth [54, 40, 45, 43]. According to this law, the expected number of distinct types in the first $t$ tokens of a sequence of words $(X_{t})_{t\in\mathbb{Z}}$ increases like a sublinear power-law function,
$V(\mathcal{X}_{1}^{t})\stackrel{{\scriptstyle*}}{{=}}t^{\beta},\quad 0<\beta<1.$ (5)
Heaps‘ law is widely viewed as a direct consequence of Zipf‘s law with $\beta=1/\alpha$ for $X_{t}=K_{t}$ [62] but the linguistic phenomenology and the statistical theory are more complicated [52, 4, 66, 67, 39, 35, 20, 29]. Understanding precisely what form Zipf‘s law implies a particular form of Heaps‘ law is an important step in our approach.
Hilberg‘s hypothesis. A reinterpretation of Shannon‘s early findings from 1950‘s [77, 78], Hilberg‘s hypothesis, also called Hilberg‘s law, was developed around 1990‘s [47, 30, 11, 19, 23, 25]. It posits that the block entropy of the first $t$ word tokens contains a sublinear power-law component,
$H(X_{1}^{t})-hn\stackrel{{\scriptstyle*}}{{=}}t^{\beta},\quad 0<\beta<1,$ (6)
where $h:=\inf_{t}H(X_{1}^{t})/t$ is the entropy rate [18].
Empirically, Hilberg‘s hypothesis understood as condition (6) with the entropy estimated by the prediction-by-partial matching (PPM) compression algorithm [75, 75, 16], holds more universally and uniformly than Zipf‘s or Heaps‘ laws. The empirical estimate $\beta\approx 0.8$ obtained for the PPM algorithm is quite stable for gigabyte-sized corpora and does not differ significantly across typologically diverse languages, using either alphabetic or ideographic scripts [81, 82]. Thus Hilberg‘s hypothesis seems a plausible candidate for a statistical language universal [82].
Santa Fe processes. Since condition (6) does not hold for IID and finite-state sources [19, 27], Hilberg‘s hypothesis might be considered a witness to large memory and complex structure of natural texts. This view is somewhat inaccurate. Large memory does not necessitate complex structure in an intuitive sense. A simple stochastic source that satisfies condition (6) is the Santa Fe process discovered by us in August 2002, described in [24, 25], and later rediscovered by Hutter [50]. The idea of the Santa Fe process is to decompose each token $X_{t}$ as a pair of a natural number $K_{t}$ and an additional bit—which is copied from a certain sequence of bits $(Z_{k})_{k\in\mathbb{N}}$ by taking the bit at position $K_{t}$. Thus, each text token $X_{t}$ may be written as a pair
$X_{t}=(K_{t},Z_{K_{t}}),$ (7)
where $(Z_{k})_{k\in\mathbb{N}}$ is the sequence of bits, called the knowledge, and $(K_{t})_{t\in\mathbb{Z}}$ is the sequence of natural numbers, called the narration.
The terms ’’knowledge‘‘ and ’’narration‘‘ were chosen because of a semantic interpretation of Santa Fe processes, discussed in [25, 26, 27]. We may interpret that pairs $(K_{t},Z_{K_{t}})$ are statements that describe bits of sequence $(Z_{k})_{k\in\mathbb{N}}$ in an arbitrary order but in a non-contradictory way. Namely, if statements $(k,Z_{k})$ and $(k^{\prime},Z_{k^{\prime}})$ describe the same bit $(k=k^{\prime})$ then they assert the same value $(Z_{k}=Z_{k^{\prime}})$. Thus, we may interpret that sequence $(Z_{k})_{k\in\mathbb{N}}$ expresses some unbounded immutable general knowledge which is only partially accessed and described in finite texts.
[…]
Neural scaling. The universal power-law behavior of information measures, when applied to natural big data, can be further confirmed by experiments with large foundation models—language or multimodal models in particular, developed in the beginning of 2020‘s [21, 73, 13, 83, 15]. These advanced statistical models build upon deep neural networks [9], word embeddings [65], and the transformer architecture [84]. Foundation models can be regarded as a game-changer in the research of language and cognition, as they allow to test probabilistic hypotheses about human language on an unprecedented scale and detail [37, 36].
A particularly salient empirical finding is that the predictive performance of these models improves with scale in a power-law fashion. The neural scaling law characterizes how the loss function of a foundation model—typically measured by the cross entropy rate on a test dataset—decreases as training data $t$, model size $n$, and compute $c$ increase [51, 44, 46, 48, 58]. Simplifying particular empirical observations and ignoring complex interactions between $t$, $n$, and $c$, this power-law relationship can be approximated as
$h(s,t,\infty,\infty)-h$ $\stackrel{{\scriptstyle*}}{{=}}t^{-\gamma_{T}},$ (8)
$h(s,\infty,n,\infty)-h$ $\stackrel{{\scriptstyle*}}{{=}}n^{-\gamma_{N}},$ (9)
$h(s,\infty,\infty,c)-h$ $\stackrel{{\scriptstyle*}}{{=}}c^{-\gamma_{C}}$ (10)
for a fixed $s<\infty$, where
$h(s,t,n,c):=\frac{\operatorname{\mathbf{E}}(-\log Q_{tnc}(X_{t+1}^{t+s}))}{s}$ (11)
is the expected cross entropy rate of a foundation model $Q_{tnc}$ tested on data $X_{t+1}^{t+s}$ and trained on data $X_{1}^{t}$ with the amount of parameters $n$ and the amount of compute $c$. The exponents satisfy $\gamma_{i}>0$.
Empirically, it is confirmed that $\gamma_{T}>\gamma_{N}$ [51]. This case is called the overparameterization. The underparameterization is the opposite regime $\gamma_{T}<\gamma_{N}$. In this paper, we will derive the neural scaling law from Hilberg‘s hypothesis but our derivation predicts underparameterization if individual parameters have a sandwich bounded Shannon entropy.
Problems discussed in the literature.
Why should the above described power laws hold for natural data? Are these laws stylized facts or can they hold exactly? How can they and their parameters be connected? Several issues have attracted a more intense treatment so far:
  • Deriving statistical laws from other principles. The history of theoretical explanation power laws is as ancient as empirical observations of these regularities [59]. Whereas we have already mentioned the most distinct kinds of explanations of Zipf‘s law [62, 68, 80, 41], we notice a recent similar activity concerning neural scaling. There is a rapidly growing body of literature that seeks to explain the neural scaling law [60, 5, 50, 61, 74, 85, 79]. Much of this work involves sophisticated mathematical frameworks, including applications of random matrix theory and techniques from theoretical physics such as Feynman diagrams. Paper [1] has also sought to link Hilberg‘s law with resource-bounded Kolmogorov complexity [55, 56] and a general theory of intelligent agents that find themselves under pressure to memorize patterns if they are rewarded for saving time.
  • Deriving statistical laws from one another. We wonder whether deriving neural scaling from principles that ignore basic quantitative linguistics is not a theoretical overkill. In fact, our notable prior, Hutter [50] derived a version of the neural scaling law in the case of a simple memory-based classification task that involves Zipf‘s law. The underlying probabilistic source in Hutter‘s paper is the exchangeable Santa Fe process (7), earlier discovered in [24, 25]. Indeed, our own activity for a long time [23, 25, 26, 27] has been connecting various forms of Zipf‘s and Heaps‘ laws with Hilberg‘s hypothesis, in the IID or general stationary setting. In particular, analyzing these laws deductively allows to infer that Zipf‘s law (4) implies Heaps‘ law (5) and Hilberg‘s law (6) with $\beta=1/\alpha$ in more or less specific settings such as Santa Fe or strongly non-ergodic processes.
  • Explaining overparameterization. Scaling laws are frequently interpreted as revealing that foundation models are overparameterized in the sense of $\gamma_{T}>\gamma_{N}$. There are papers that seek to explain this issue [8, 6, 89, 7, 61]. Yet, in view of the results of our paper, we suppose that overparameterization may be an artifact. Namely, we suppose that larger models effectively use fewer bits of information per real parameter, blurring the operational meaning of parameter count.
  • Studying departures from simple formulas. It has been known in quantitative linguistics that statistical laws of language are stylized facts rather than follow simple exact formulas. A part of the problem of empirical studies is that expectations are easier to investigate than variances, whereas variances are large for language data, due to Taylor‘s law [53, 82] and burstiness of words [3]. Known phenomena that break a simple picture of power laws in language include two-regime rank-frequency plots [34, 69, 72, 39, 86], log-log convexity of the vocabulary growth [35], and monotone decreasing or $U$-shaped hapax rates [32, 29]. What is fascinating, the empirical marginal distribution of words can be still quite well modeled by non-parametric and parametric urn models—IID sources of words [4, 66, 67, 20, 29].
Guided by the successes of parametric IID models in simultaneous modeling of several quantitative laws such as Zipf‘s law, Heaps‘ law, and the hapax rate [4, 66, 67, 20, 29], we envision that a similarly systematic approach to Hilberg‘s law and neural scaling may succeed as well. A good theory of language and foundation models should predict the functional forms of all these laws simultaneously and predict the values of their parameters. In this paper, we want to supply some basic connections among four laws in a relatively general setting.
Overview of our results.
We will systematically prove connections among four statistical laws in particular settings that have not been discussed so far. The present paper develops and strengthens the ideas from an earlier and unpublished attempt to attack this topic [28]. We would like to supply a simple-minded baseline that, in contrast to [28], takes into account also the effect of limited amount of training compute $c$. The main contributions of the present paper are organized as a chain of following derivations:
Zipf‘s law $\stackrel{{\scriptstyle(A)}}{{\Rightarrow}}$ Heaps‘ law $\stackrel{{\scriptstyle(B)}}{{\Rightarrow}}$ Hilberg‘s law $\stackrel{{\scriptstyle(C)}}{{\Rightarrow}}$ neural scaling.
Each of the above implications involves different assumptions but we strive at the most general settings such as non-stationary processes. We remark that the conditions for implication $(C)$ (non-stationarity) are more general than that for result $(B)$ (stationarity and Santa Fe decomposition), and those are more general than the requirements for implication $(A)$ (stationarity and mixing). We also notice that the derivation of the neural scaling law rests on the differential Heaps and Hilberg laws (15)–(16), which strengthen the plain Heaps and Hilberg laws (5)–(6). Noticing a practical need for such differential laws seems a novel idea of this paper.
Let us begin with result $(C)$, to be followed by $(B)$ and $(A)$. Let $Q_{tnc}$ be a random probability measure that satisfies
$H(Q_{tnc}|X_{1}^{t})$ $\stackrel{{\scriptstyle*}}{{<}}c,$ (12)
$H(Q_{tnc})$ $\stackrel{{\scriptstyle*}}{{<}}n.$ (13)
Bounds (12)–(13) model the constraints on the amounts of training data $t$, parameters $n$, and compute $c$. (The validity of these assumptions is discussed in the next paragraph.) In the result $(C)$, for a sufficiently small test sample length $s$, we derive the neural scaling law of form
$\sup_{k\geq t}\frac{\operatorname{\mathbf{E}}(-\log Q_{tnc}(X_{k+1}^{k+s}))}{s}-h$
$\stackrel{{\scriptstyle*}}{{>}}\max\left\{t^{\beta-1}\left(\frac{1-\sqrt{\frac{ct^{-\beta}}{1-\beta}}}{1+\sqrt{\frac{ct^{-\beta}}{1-\beta}}}\right)^{1-\beta}-\frac{c}{t}\left(\frac{1-\sqrt{\frac{ct^{-\beta}}{1-\beta}}}{2\sqrt{\frac{ct^{-\beta}}{1-\beta}}}\right),\frac{2^{\beta}-1+\beta}{2}\left(\frac{n}{1-\beta}\right)^{1-\frac{1}{\beta}}\right\}.$ (14)
Result $(C)$ requires an arbitrary (non-stationary) process $(X_{t})_{t\in\mathbb{Z}}$ and a differential form of Hilberg‘s law
$\sup_{k\geq t}\frac{H(X_{k+1}^{k+s}|X_{1}^{t})}{s}-h$ $\stackrel{{\scriptstyle*}}{{>}}(t+s)^{\beta-1}.$ (15)
To match this result, we derive implication $(B)$ for an arbitrary stationary Santa Fe process (7) with $H(K_{t})<\infty$ and an analogous differential form of Heaps‘ law for narration
$\sup_{k\geq t}\frac{V(\mathcal{K}_{k+1}^{k+s}\setminus\mathcal{K}_{1}^{t})}{s}$ $\stackrel{{\scriptstyle*}}{{>}}(t+s)^{\beta-1}.$ (16)
Finally, we show that this form Heaps‘ law is satisfied by sufficiently strongly mixing narrations with sufficiently heavy-tailed marginal distributions. In particular, we derive result $(A)$ for a stationary process $(K_{t})_{t\in\mathbb{Z}}$ with a mixing condition $\Psi<\infty$, where
$\Psi$ $:=\sup_{t,k}\frac{\sum_{i=1}^{t}\sum_{j=1}^{t}P(K_{i}=k,K_{j}=k)-tP(K_{0}=k)}{t(t-1)(P(K_{0}=k))^{2}},$ (17)
and an approximate Zipf law
$P(K_{0}=k)$ $\stackrel{{\scriptstyle*}}{{=}}k^{-1/\beta}.$ (18)
Open problems.
Our analysis highlights some gaps that need to be addressed in the future research:
  • Modeling bounds on compute. Bounds (12)–(13) are formulated in a purely information-theoretic sense. We may do it because the entropy of a discrete object cannot be essentially larger than the description length of this object. Thus limiting the amount of a certain resource implies constraining the respective entropy. In particular, we assume that model $Q_{tnc}$ may be a stochastic function of training data $X_{1}^{t}$. We suppose that training may involve randomization in the amount proportional to the amount of the training compute. If there is no randomization involved, constraint (12) should be formulated in terms of a resource-bounded Kolmogorov complexity [55, 56] rather than entropy. Effectively, constraining the compute bounds the amount of time that we have to infer model $Q_{tnc}$ from training data $X_{1}^{t}$. Some ideas of paper [1] may be useful in this approach.
  • Overparameterization. The result stated in equation (14) suggests theoretical optimality of underparameterization $\gamma_{T}\leq\gamma_{N}$, contrary to the empirical optimality of overparameterization $\gamma_{T}>\gamma_{N}$ [51]. Indeed, suppose that (14) holds. Then we derive
    $\gamma_{T}$ $\leq 1-\beta,$ (19)
    $\gamma_{N}$ $\leq\frac{1}{\beta}-1.$ (20)
    Thus understanding why overparameterized models perform empirically better requires an explanation. One such explanation may be that condition (13) does not capture the fact that a single real-valued parameter (a single weight) may carry potentially an infinite amount of information (theoretically, it is an infinite sequence of bits). Thus, the effective complexity of the model is greater than the number of real-valued parameters. The empirical advantage of overparameterization may mean that larger models effectively use fewer digits of the binary expansion per real-valued parameter. If this is true then the established notion of overparameterization may be misleading and quantifying the true degrees of freedom remains an open question.
  • Arbitrariness of word shapes. Having a general implication from Hilberg‘s law to neural scaling, the burden of theoretical explanations shifts to stating why Hilberg‘s hypothesis may be sound. In general, Hilberg‘s hypothesis can be viewed as a sort of Zipf‘s law for tokens that are internally random enough like in the Santa Fe process (7). The internal randomness of these tokens can be potentially connected to semantic or linguistic considerations [25, 26, 64]. In particular, we may expect that Hilberg‘s law arises if we have Zipf‘s law for words that have sufficiently arbitrary shapes, cf. [22]. Proving such a claim formally requires a longer excursion to universal coding and involves results that exceed those of papers [25, 26, 27]. For this reason, we postpone this theoretical development to another article.
Organization of the article.
The organization of this paper is as follows. Section 2 introduces notation and preliminaries. In particular, Section 2.1 treats fundamental inequalities for Shannon entropy and expected cardinality. Section 2.2 discusses the rates of entropy and expected cardinality for arbitrary (non-stationary) processes. Section 2.3 handles the block entropy and the expected block cardinality for stationary processes. Section 2.4 analyzes the vocabulary spectrum elements for IID processes. Section 3 develops the main chain of implications. In particular, Section 3.1 presents the derivation of Heaps‘ law from Zipf‘s law. Section 3.2 develops the implication from Heaps‘ law to Hilberg‘s hypothesis. Section 3.3 establishes the derivation of neural scaling from Hilberg‘s hypothesis. Section 4 concludes the article.
2 Preliminaries
2.1 Fundamental inequalities
We will be developing parallel results for the Shannon entropy and the expected number of types. Our reasoning is based on simple but systematic information-theoretic considerations and the general spirit of the $I$-measure [49]. For the textbook treatment and the background, we refer to [18, 88]. A handy tool is also a generalization of Shannon information measures to arbitrary $\sigma$-fields [26, Sections 5.3 and 5.4].
In general, both the Shannon entropy and the expected cardinality are subadditive and enjoy the triangle inequality.
Proposition 1 (subadditivity).
For random variables $X,Y,Q$, we have
$H(X,Y|Q)$ $\leq H(X|Q)+H(Y|Q).$ (21)
Proof.
The claim follows by the chain rule
$H(X,Y|Q)$ $=H(X|Q)+H(Y|Q,X)$ (22)
and inequality $H(Y|Q,X)\leq H(Y|Q)$. ∎
Proposition 2 (subadditivity).
For random sets $\mathcal{X},\mathcal{Y},\mathcal{Q}$, we have
$V(\mathcal{X}\cup\mathcal{Y}\setminus\mathcal{Q})$ $\leq V(\mathcal{X}\setminus\mathcal{Q})+V(\mathcal{Y}\setminus\mathcal{Q}).$ (23)
Proof.
The claim follows by the chain rule
$V(\mathcal{X}\cup\mathcal{Y}\setminus\mathcal{Q})$ $=V(\mathcal{X}\setminus\mathcal{Q})+V(\mathcal{Y}\setminus\mathcal{Q}\cup\mathcal{X})$ (24)
and inequality $V(\mathcal{Y}\setminus\mathcal{Q}\cup\mathcal{X})\leq V(\mathcal{Y}\setminus\mathcal{Q})$. ∎
Proposition 3 (triangle inequality).
For random variables $X,Y,Q$, we have
$H(X|Y)$ $\leq H(X|Q)+H(Q|Y).$ (25)
Proof.
The claim follows by the chain rule
$H(X,Q|Y)$ $=H(X|Y,Q)+H(Q|Y)$ (26)
and inequalities $H(X|Y)\leq H(X,Q|Y)$ and $H(X|Y,Q)\leq H(X|Q)$. ∎
Proposition 4 (triangle inequality).
For random sets $\mathcal{X},\mathcal{Y},\mathcal{Q}$, we have
$V(\mathcal{X}\setminus\mathcal{Y})$ $\leq V(\mathcal{X}\setminus\mathcal{Q})+V(\mathcal{Q}\setminus\mathcal{Y}).$ (27)
Proof.
The claim follows by the chain rule
$V(\mathcal{X}\cup\mathcal{Q}\setminus\mathcal{Y})$ $=V(\mathcal{X}\setminus\mathcal{Y}\cup\mathcal{Q})+V(\mathcal{Q}\setminus\mathcal{Y})$ (28)
and inequalities $V(\mathcal{X}\setminus\mathcal{Y})\leq V(\mathcal{X}\cup\mathcal{Q}\setminus\mathcal{Y})$ and $V(\mathcal{X}\setminus\mathcal{Y}\cup\mathcal{Q})\leq V(\mathcal{X}\setminus\mathcal{Q})$. ∎
We note that following these analogies, the counterpart of mutual information I(X;Y):=H(X)+H(Y)H(X,Y)I(X;Y):=H(X)+H(Y)-H(X,Y) is the expected cardinality of intersection $V(\mathcal{X}\cap\mathcal{Y})=V(\mathcal{X})+V(\mathcal{Y})-V(\mathcal{X}\cup\mathcal{Y})$, cf. [49]. However, this is an incomplete analogy since in general there is no random variable $Q=f(X,Y)$ such that $H(Q)=I(X;Y)$, cf. [38, 87].
There is also an important bridging inequality for cross entropy of random measures.
Proposition 5 (source coding).
For a random probability measure $Q$ applied to another random variable $X$, we have
$\operatorname{\mathbf{E}}(-\log Q(X))$ $\geq H(X|Q).$ (29)
Proof.
We have
$\operatorname{\mathbf{E}}\left(-\log Q(X)\right)$ $=\operatorname{\mathbf{E}}\operatorname{\mathbf{E}}\left(-\log Q(X)\middle|Q\right)$
$=\operatorname{\mathbf{E}}\operatorname{\mathbf{E}}\left(-\log P(X|Q)\middle|Q\right)+\operatorname{\mathbf{E}}\operatorname{\mathbf{E}}\left(\log\frac{P(X|Q)}{Q(X)}\middle|Q\right)$
$\geq\operatorname{\mathbf{E}}\operatorname{\mathbf{E}}\left(-\log P(X|Q)\middle|Q\right)$
$=\operatorname{\mathbf{E}}\left(-\log P(X|Q)\right)$
$=H(X|Q)$
$\geq H(X)-H(Q),$ (30)
where we have used the law of total expectation and non-negativity of the Kullback-Leibler divergence. ∎
2.2 Arbitrary processes
Let us inspect the rates of the Shannon entropy and the expected cardinality for general stochastic processes (over a countable alphabet). We may define theses rates as follows.
Definition 1.
Let $(X_{t})_{t\in\mathbb{Z}}$ be an arbitrary stochastic process. We define the entropy rate
$h$ $:=\inf_{s\in\mathbb{N}}\sup_{k\geq 0}\frac{H(X_{k+1}^{k+s})}{s}.$ (31)
Definition 2.
Let $(K_{t})_{t\in\mathbb{Z}}$ be an arbitrary stochastic process. We define the expected cardinality rate
$v$ $:=\inf_{s\in\mathbb{N}}\sup_{k\geq 0}\frac{V(\mathcal{K}_{k+1}^{k+s})}{s}.$ (32)
Having subadditivity (21) and (23), we can show that these rates can be equivalently expressed somewhat differently.
Proposition 6.
Let $(X_{t})_{t\in\mathbb{Z}}$ be a stochastic process such that $h<\infty$ for the entropy rate (31). For an arbitrary $t\geq 0$, we have
$h=\lim_{s\to\infty}\sup_{k\geq t}\frac{H(X_{k+1}^{k+s}|X_{1}^{t})}{s}=\inf_{s\in\mathbb{N}}\sup_{k\geq t}\frac{H(X_{k+1}^{k+s}|X_{1}^{t})}{s}.$ (33)
Remark: Hence $\sup_{k\geq t}H(X_{k+1}^{k+s}|X_{1}^{t})\geq hs$.
Proof.
By inequality (21), we notice subadditivity
$\sup_{k\geq t}H(X_{k+1}^{k+s+r}|X_{1}^{t})$ $\leq\sup_{k\geq t}\left[H(X_{k+1}^{k+s}|X_{1}^{t})+H(X_{k+s+1}^{k+s+r}|X_{1}^{t})\right]$
$\leq\sup_{k\geq t}H(X_{k+1}^{k+s}|X_{1}^{t})+\sup_{k\geq t}H(X_{k+1}^{k+r}|X_{1}^{t}).$ (34)
Hence by the Fekete lemma [33], we obtain
$h(t):=\lim_{s\to\infty}\sup_{k\geq t}\frac{H(X_{k+1}^{k+s}|X_{1}^{t})}{s}=\inf_{s\in\mathbb{N}}\sup_{k\geq t}\frac{H(X_{k+1}^{k+s}|X_{1}^{t})}{s}.$ (35)
It suffices to show $h(t)=h$. Since $h<\infty$, we have $\sup_{t\geq 0}H(X_{t})<\infty$. Hence, by the chain rule and a simple calculation, we obtain a uniform bound
$\left|H(X_{k+1}^{k+s})-H(X_{k+t+1}^{k+t+s})\right|\leq B(t)<\infty.$ (36)
For the same reason, we also have
$\left|H(X_{k+t+1}^{k+t+s})-H(X_{k+t+1}^{k+t+s}|X_{1}^{t})\right|\leq D(t)<\infty.$ (37)
Chaining these two sandwich bounds and taking the supremums over $k$ and infimums over $s$, we infer $h(t)=h$. ∎
Proposition 7.
Let $(K_{t})_{t\in\mathbb{Z}}$ be an arbitrary stochastic process. For an arbitrary $t\geq 0$, we have
$v=\lim_{s\to\infty}\sup_{k\geq t}\frac{V(\mathcal{K}_{k+1}^{k+s}\setminus\mathcal{K}_{1}^{t})}{s}=\inf_{s\in\mathbb{N}}\sup_{k\geq t}\frac{V(\mathcal{K}_{k+1}^{k+s}\setminus\mathcal{K}_{1}^{t})}{s}.$ (38)
Remark: Hence $\sup_{k\geq t}V(\mathcal{K}_{k+1}^{k+s}\setminus\mathcal{K}_{1}^{t})\geq vs$.
Proof.
Mutatis mutandis, the same as the proof of Proposition 6. ∎
2.3 Stationary processes
For stationary processes $(X_{t})_{t\in\mathbb{Z}}$ and $(K_{t})_{t\in\mathbb{Z}}$, we denote the the block Shannon entropy and the expected number of types
$H(t)$ $:=H(X_{1}^{t}),$ (39)
$V(t)$ $:=V(\mathcal{K}_{1}^{t}).$ (40)
Let us define the finite difference operator $\Delta f(t):=f(t+1)-f(t)$. Function $f(t)$ is called positive, growing, and concave if $f(t)\geq 0$, $\Delta f(t)\geq 0$, and $\Delta^{2}f(t)\leq 0$, respectively.
We have two analogous statements. We recall the results for entropy to apply them by analogy to the expected cardinality.
Proposition 8.
Let a stationary process $(X_{t})_{t\in\mathbb{Z}}$. For $t\in\mathbb{N}$, we claim that:
  1. 1.
    Function $t\mapsto H(t)$ is positive, growing, and concave and $H(0)=0$.
  2. 2.
    Function $t\mapsto H(t)/t$ is decreasing.
  3. 3.
    We have $h=\lim_{t\to\infty}H(t)/t\geq 0$.
Proof.
See [26, §5.2]. In general, we have
$H(t)$ $=H(X_{1}^{t})\geq 0,$ (41)
$\Delta H(t)$ $=H(X_{t+1}|X_{1}^{t})\geq 0,$ (42)
$\Delta^{2}H(t)$ $=-I(X_{0};X_{t+1}|X_{1}^{t})\leq 0.$ (43)
Function $t\mapsto\Delta H(t)$ is decreasing since $\Delta^{2}H(t)\leq 0$. Hence
$\frac{H(t+1)}{t+1}=\frac{\sum_{i=0}^{t}\Delta H(i)}{t+1}$ $\leq\frac{\sum_{i=0}^{t-1}\Delta H(i)+\frac{\sum_{i=0}^{t-1}\Delta H(i)}{t}}{t+1}$
$=\frac{\sum_{i=0}^{t-1}\Delta H(i)}{t}=\frac{H(t)}{t}.$ (44)
Thus function $t\mapsto H(t)/t$ is decreasing and limit $\lim_{t\to\infty}H(t)/t$ exists. It equals $h$ defined in (31) by stationarity. ∎
Proposition 9.
Let a stationary process $(K_{t})_{t\in\mathbb{Z}}$. For $t\in\mathbb{N}$, we claim that:
  1. 1.
    Function $t\mapsto V(t)$ is positive, growing, and concave and $V(0)=0$.
  2. 2.
    Function $t\mapsto V(t)/t$ is decreasing and $V(t)/t\leq 1$.
  3. 3.
    We have $v=\lim_{t\to\infty}V(t)/t=0$.
Proof.
The proof is analogous to the proof of Proposition 8 except for the last claim. In particular, we have
$V(t)$ $=V(\mathcal{K}_{1}^{t})\geq 0,$ (45)
$\Delta V(t)$ $=V(\mathcal{K}_{t+1}\setminus\mathcal{K}_{1}^{t})\geq 0,$ (46)
$\Delta^{2}V(t)$ $=-V(\mathcal{K}_{0}\cap\mathcal{K}_{t+1}\setminus\mathcal{K}_{1}^{t})\leq 0.$ (47)
The proof of $\lim_{t\to\infty}V(t)/t=0$ is as follows. Without loss of generality, we assume that the alphabet is the set of natural numbers. Let $g(t)$ be an arbitrary function. Generalizing an idea used by Khmaladze [52] for IID processes, we observe
$V(t)=\operatorname{\mathbf{E}}\sum_{k=1}^{\infty}{\bf 1}{\left\{k\in\mathcal{K}_{1}^{t}\right\}}=\sum_{k=1}^{\infty}P(k\in\mathcal{K}_{1}^{t})\leq g(t)+\sum_{k>g(t)}P(k\in\mathcal{K}_{1}^{t}),$ (48)
whereas the union bound and stationarity yield
$P(k\in\mathcal{K}_{1}^{t})=P(K_{1}=k\lor\ldots\lor K_{t}=k)\leq tP(K_{0}=k).$ (49)
Thus we have bound
$V(t)\leq g(t)+tP(K_{0}>g(t))$ (50)
that holds for any function $g(t)$. In particular, for an $\epsilon>0$, we may take $g(t)=\epsilon t/2$. For all sufficiently large $t$, we observe $P(K_{0}>g(t))\leq\epsilon/2$. Hence, for these $t$, we have $V(t)/t\leq g(t)/t+P(K_{0}>g(t))\leq\epsilon$. By arbitrariness of $\epsilon$, we derive $\lim_{t\to\infty}V(t)/t=0$. ∎
Besides the expected number of all types $V(t)$, let us introduce expected spectrum elements $V(t|m)$, defined as the expected number of types that occur exactly $m$ times.
Definition 3.
The expected spectrum elements are defined as
$V(t|m):=\sum_{k}P\left(\sum_{i=1}^{t}{\bf 1}{\left\{K_{i}=k\right\}}=m\right).$ (51)
In particular, $V(t|1)$ is the expected number of types that occur once. These are called hapax legomena in Greek or, succinctly, hapaxes in English. There is a general upper bound for the expected number of hapaxes in the stationary case.
Proposition 10.
For a stationary process $(K_{t})_{t\in\mathbb{Z}}$, we have
$\frac{V(t|1)}{t}\leq\Delta V\left(\left\lceil\frac{t}{2}\right\rceil\right).$ (52)
Proof.
For a stationary process, we observe that
$V(t|1)$ $=\sum_{i=1}^{t}P\left(K_{i}\not\in\mathcal{K}_{1}^{i-1}\cup\mathcal{K}_{i+1}^{t}\right)$
$\leq\sum_{i=1}^{t}\min\left\{P\left(K_{i}\not\in\mathcal{K}_{1}^{i-1}\right),P\left(K_{i}\not\in\mathcal{K}_{i+1}^{t}\right)\right\}$
$=\sum_{i=1}^{t}\min\left\{\Delta V(i-1),\Delta V(t-i)\right\}\leq t\Delta V\left(\left\lceil\frac{t}{2}\right\rceil\right)$ (53)
since $t\mapsto\Delta V(t)$ is decreasing. ∎
2.4 IID processes
So far, the statements for Shannon entropy and expected cardinality were mirror-like. However, for more specific processes, the analogy between these two functionals is rather as follows: The expected cardinality applies to IID processes as the Shannon entropy applies to exchangeable Santa Fe processes. This analogy will generalized formally in Section 3.2. Now let us comment on the expected cardinality for IID processes, as it is simpler to analyze. We notice that the expected cardinality for IID sources enjoys further properties: It is a Hausdorff sequence—a discrete-time analog of the Bernstein function.
Combining creatively several sources such as [4, 2, 76], the development is as follows. By the Bernoulli scheme, for an IID process $(K_{t})_{t\in\mathbb{Z}}$, we have
$V(t)$ $=\sum_{k}(1-(1-p_{k})^{t}),$ (54)
$V(t|m)$ $=\begin{cases}\sum_{k}\binom{t}{m}p_{k}^{m}(1-p_{k})^{t-m},&1\leq m\leq t,\\ 0,&m>t,\end{cases}$ (55)
where $p_{k}:=P(K_{t}=k)$. We notice consistency conditions
$\sum_{m=1}^{\infty}V(t|m)$ $=V(t),$ $\sum_{m=1}^{\infty}mV(t|m)$ $=t.$ (56)
Moreover, since $\Delta(1-p_{k})^{t}=-p_{k}(1-p_{k})^{t}$, we also observe that the spectrum elements can be computed from the expected number of types by taking differences,
$V(t|m)$ $=\begin{cases}(-1)^{m+1}\binom{t}{m}\Delta^{m}V(t-m),&1\leq m\leq t,\\ 0,&m>t.\end{cases}$ (57)
Expression (57) reminds of a known concept.
Definition 4 (cf. [2]).
A sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ is called a Hausdorff sequence if $v(t)\geq 0$ and $(-1)^{m+1}\Delta^{m}v(t)\geq 0$ for all $m\in\mathbb{N}$ and $n\in\mathbb{N}\cup\left\{0\right\}$. A sequence $u:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ is called completely alternating if $u(t)=\Delta v(t)$ for a certain Hausdorff sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$. We call a Hausdorff sequence $v(t)$ standard if $v(0)=0$ and $\Delta v(0)=1$.
The name Hausdorff sequence is non-standard itself. We have coined it by an analogy to the standard term Bernstein function, which is the continuous time-analog, applying derivatives rather than differences [76].
Any Hausdorff sequence can be expressed as a convex combination of sequences $t\mapsto p^{-1}\left(1-(1-p)^{t}\right)$ for varying $p>0$. This result is known as the Hausdorff moment theorem [42]. It is a discrete-time analog of the Bernstein theorem on completely monotone functions [10], also known as the Lévy-Khintchine representation in probability [57].
Proposition 11 ([2]).
A sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ is a Hausdorff sequence if and only if there exists a unique non-negative measure $F$ on $[0,1]$ such that
$v(t)$ $=tF(\left\{0\right\})+\int_{(0,1)}\frac{1-(1-p)^{t}}{p}dF(p)+F(\left\{1\right\}),$ (58)
$\Delta v(t)$ $=F(\left\{0\right\})+\int_{(0,1)}(1-p)^{t}dF(p).$ (59)
Additionally, we see that a Hausdorff sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ is standard if and only if $F(\left\{1\right\})=0$ and the measure $F$ is a probability measure, namely,
$F(\left\{0\right\})+\int_{(0,1)}dF(p)=1.$ (60)
We call measure $F$ the Hausdorff measure.
We notice that the Newton formula $(1-\Delta)^{r}=\sum_{m=0}^{r}\binom{r}{m}(-\Delta)^{m}$ can be rewritten in the following form, which resembles the Taylor expansion,
$v(t-r)$ $=\sum_{m=0}^{r}(-1)^{m}\binom{r}{m}\Delta^{m}v(t-m).$ (61)
By formula (61), consistency conditions analogous to (56) can be verified easily. Namely, a Hausdorff sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ with spectrum elements
$v(t|m)$ $:=\begin{cases}(-1)^{m+1}\binom{t}{m}\Delta^{m}v(t-m),&1\leq m\leq t,\\ 0,&m>t.\end{cases}$ (62)
satisfies conditions $\sum_{m=1}^{t}v(t|m)=v(t)$ and $\sum_{m=1}^{t}mv(t|m)=t$ if and only if it is standard.
To complete this discussion, the expected number of types $V(t)$ is a standard Hausdorff sequence and it has a Hausdorff probability measure supported on discrete points, namely,
$F(A)=\sum_{k}p_{k}{\bf 1}{\left\{p_{k}\in A\right\}}.$ (63)
In Section 3.1, we approximate this measure by a non-atomic measure so as to derive two symmetric bounds for the conditional Heaps law from a bound for Zipf‘s law.
3 Implications
3.1 Zipf‘s law implies Heaps‘ law
Let us recall from Section 2.4 that the expected number of hapaxes for IID processes satisfies
$V(t+1|1)=(t+1)\Delta V(t).$ (64)
The growth rate of the number of hapaxes $V(t+1|1)$ is controlled by the tail of the marginal distribution. In fact, there are two simple symmetric conditions on the tail of the marginal distribution that sandwich bound the number of hapaxes and lead to a differential Heaps law.
We continue notation $p_{k}:=P(K_{t}=k)$ in the following.
Proposition 12.
For an IID process $(K_{t})_{t\in\mathbb{Z}}$, suppose that
$\sum_{k}{\bf 1}{\left\{p_{k}\geq p\right\}}\leq C_{1}p^{-\beta}$ (65)
for some $\beta\in(0,1)$, $C_{1}>0$, and all $p>0$. Then
$\Delta V(t)\leq C_{2}t^{\beta-1},$ (66)
where $C_{2}:=\Gamma(1-\beta)C_{1}$.
Proof.
We may bound
$\Delta V(t)$ $=\frac{V(t+1|1)}{t+1}=\sum_{k}p_{k}(1-p_{k})^{t}$
$\leq\sum_{k}\int_{0}^{p_{k}}(1-p)^{t}dp=\int_{0}^{1}\left(\sum_{k}{\bf 1}{\left\{p_{k}\geq p\right\}}\right)(1-p)^{t}dp$
$\leq C_{1}\int_{0}^{1}p^{-\beta}(1-p)^{t}dp.$ (67)
Further evaluation yields
$\int_{0}^{1}(1-p)^{t}p^{-\beta}dp$ $=\frac{\Gamma(t+1)\Gamma(1-\beta)}{\Gamma(t+2-\beta)}\leq\Gamma(1-\beta)t^{\beta-1}.$ (68)
The reverse bound can be demonstrated for a larger class of processes. For this aim, let us define a suitable mixing coefficient.
Definition 5.
For a stationary process $(K_{t})_{t\in\mathbb{Z}}$, denote the mixing coefficient
$\Psi:=\sup_{t,k}\frac{\sum_{i=1}^{t}\sum_{j=1}^{t}P(K_{i}=k,K_{j}=k)-tp_{k}}{t(t-1)p_{k}^{2}}.$ (69)
In particular, for an IID process, we have $\Psi=1$. For a stationary ergodic process, we have $\lim_{t\to\infty}t^{-1}\sum_{i=1}^{t}P(K_{0}=k,K_{i}=k)=p_{k}^{2}$ [26, Theorem 4.9]. For a comprehensive survey of various mixing coefficients and mixing conditions, we refer to [12]. For our purposes, it is sufficient to assume that $\Psi$ is finite.
Proposition 13.
For a stationary process $(K_{t})_{t\in\mathbb{Z}}$ with $\Psi<\infty$, suppose that
$\sum_{k}p_{k}{\bf 1}{\left\{p_{k}\leq p\right\}}\geq C_{3}p^{1-\beta}$ (70)
holds for some $\beta\in(0,1)$, $C_{3}>0$, and all $p>0$. Then we have
$\Delta V(t)\geq C_{4}t^{\beta-1},$ (71)
where $C_{4}:=(4\Psi)^{\beta-1}C_{3}/2$.
Proof.
By (52), using the union bound, we may write
$\Delta V(t)$ $\geq\frac{V(2t|1)}{2t}=\frac{1}{2t}\sum_{i=1}^{2t}P\left(K_{i}\not\in\mathcal{K}_{1}^{i-1}\cup\mathcal{K}_{i+1}^{2t}\right)$
$=\sum_{k}\frac{1}{2t}\sum_{i=1}^{2t}P\left(K_{i}=k\not\in\mathcal{K}_{1}^{i-1}\cup\mathcal{K}_{i+1}^{2t}\right)$
$=\sum_{k}\max\left\{0,\frac{1}{2t}\sum_{i=1}^{2t}P\left(K_{i}=k\not\in\mathcal{K}_{1}^{i-1}\cup\mathcal{K}_{i+1}^{2t}\right)\right\}$
$\geq\sum_{k}\max\left\{0,2p_{k}-\frac{1}{2t}\sum_{i=1}^{2t}\sum_{j=1}^{2t}P(K_{i}=k,K_{j}=k)\right\}$
$\geq\sum_{k}p_{k}\max\left\{0,1-\Psi(2t-1)p_{k}\right\}\geq\sum_{k}p_{k}\max\left\{0,1-2\Psi tp_{k}\right\}$
$\geq\frac{1}{2}\sum_{k}p_{k}{\bf 1}{\left\{p_{k}\leq\frac{1}{4\Psi t}\right\}}\geq\frac{C_{3}(4\Psi t)^{\beta-1}}{2}.$ (72)
Now we can show that conditions (65) and (70) hold under a broadly understood Zipf law.
Example 1.
Consider an approximate Zipf law
$C_{5}k^{-\alpha}\leq p_{k}\leq C_{6}k^{-\alpha}$ (73)
for some $\alpha>1$, $C_{5},C_{6}>0$, and all $k\in\mathbb{N}$. Then for $\beta=1/\alpha$, we have
$\sum_{k}{\bf 1}{\left\{p_{k}\geq p\right\}}$ $\leq\int_{0}^{\infty}{\bf 1}{\left\{C_{6}k^{-\alpha}\geq p\right\}}dk$
$=\int_{0}^{\infty}{\bf 1}{\left\{k\leq\left(\frac{p}{C_{6}}\right)^{-\beta}\right\}}dk=\left(\frac{p}{C_{6}}\right)^{-\beta},$ (74)
$\sum_{k}p_{k}{\bf 1}{\left\{p_{k}\leq p\right\}}$ $\geq\int_{1}^{\infty}C_{5}k^{-\alpha}{\bf 1}{\left\{C_{6}k^{-\alpha}\leq p\right\}}dk$
$=\int_{1}^{\infty}C_{5}k^{-\alpha}{\bf 1}{\left\{k\geq\left(\frac{p}{C_{6}}\right)^{-\beta}\right\}}dk=\frac{C_{5}}{1/\beta-1}\left(\frac{p}{C_{6}}\right)^{1-\beta}.$ (75)
Thus conditions (65) and (70) hold for some constants $C_{1},C_{3}>0$. In consequence, an approximate Zipf law (73) implies the differential Heaps law (66) if the process is IID and it implies the differential Heaps law (71) if the process is stationary and $\Psi<\infty$.
3.2 Heaps‘ law implies Hilberg‘s law
Now we will formalize the remark made in Section 2.4. Namely, we will use the correspondence that the expected cardinality applies to arbitrary (stationary) processes as the Shannon entropy applies to (stationary) Santa Fe processes. In particular, the differential Heaps law (71) for a stationary narration implies a differential Hilberg law.
Proposition 14.
Consider a Santa Fe process $(X_{t})_{t\in\mathbb{Z}}$ such that decomposition (7) holds, where narration $(K_{t})_{t\in\mathbb{Z}}$ is an arbitrary process over natural numbers and knowledge $(Z_{k})_{k\in\mathbb{N}}$ is a sequence of independent random variables with $H(Z_{k})\in[C_{7},C_{8}]\subset(0,\infty)$ independent of narration $(K_{t})_{t\in\mathbb{Z}}$. The following assertions are true:
  1. 1.
    If $v=0$ for process $(K_{t})_{t\in\mathbb{Z}}$ then
    $\sup_{k\geq t}\frac{H(X_{k+1}^{k+s}|X_{1}^{t})}{s}-h$ $\geq C_{7}\inf_{k\geq t}\frac{V(\mathcal{K}_{k+1}^{k+s}\setminus\mathcal{K}_{1}^{t})}{s}.$ (76)
  2. 2.
    If process $(K_{t})_{t\in\mathbb{Z}}$ is stationary and $H(K_{t})<\infty$ then
    $\sup_{k\geq t}\frac{H(X_{k+1}^{k+s}|X_{1}^{t})}{s}-h$ $\geq C_{7}\sup_{k\geq t}\frac{V(\mathcal{K}_{k+1}^{k+s}\setminus\mathcal{K}_{1}^{t})}{s}\geq C_{7}\Delta V(t+s).$ (77)
Proof.
We may decompose
$H(X_{1}^{t})$ $=H(K_{1}^{t})+H(X_{1}^{t}|K_{1}^{t})$
$=H(K_{1}^{t})+H(\left\{(k,Z_{k}):k\in\mathcal{K}_{1}^{t}\right\}|\mathcal{K}_{1}^{t})$
$=H(K_{1}^{t})+V(\mathcal{K}_{1}^{t}).$ (78)
Similarly, after a longer calculation, we obtain
$H(X_{k+1}^{k+s}|X_{1}^{t})=H(K_{k+1}^{k+s}|K_{1}^{t})+V(\mathcal{K}_{k+1}^{k+s}\setminus\mathcal{K}_{1}^{t}).$ (79)
Therefore, we may bound
$\sup_{k\geq t}\frac{H(K_{k+1}^{k+s}|K_{1}^{t})}{s}$ $\leq\sup_{k\geq t}\frac{H(X_{k+1}^{k+s}|X_{1}^{t})}{s}$
$\leq\sup_{k\geq t}\frac{H(K_{k+1}^{k+s}|K_{1}^{t})}{s}+C_{8}\sup_{k\geq t}\frac{V(\mathcal{K}_{k+1}^{k+s}\setminus\mathcal{K}_{1}^{t})}{s}$ (80)
Taking $s\to\infty$ and using $v=0$, we obtain that the generalized entropy rates of processes $(K_{t})_{t\in\mathbb{Z}}$ and $(X_{t})_{t\in\mathbb{Z}}$ are equal. Hence
$\sup_{k\geq t}\frac{H(X_{k+1}^{k+s}|X_{1}^{t})}{s}$ $\geq\sup_{k\geq t}\frac{H(K_{k+1}^{k+s}|K_{1}^{t})}{s}+C_{7}\inf_{k\geq t}\frac{V(\mathcal{K}_{k+1}^{k+s}\setminus\mathcal{K}_{1}^{t})}{s}$
$\geq h+C_{7}\inf_{k\geq t}\frac{V(\mathcal{K}_{k+1}^{k+s}\setminus\mathcal{K}_{1}^{t})}{s}.$ (81)
Regrouping yields claim (76).
As for the second part, we observe that condition
$\lim_{k\to\infty}H(K_{t+1}|K_{-k}^{t})=H(K_{t+1}|K_{-\infty}^{t})$ (82)
holds in particular if the alphabet of process $(K_{t})_{t\in\mathbb{Z}}$ is finite (see [26, Sections 5.3 and 5.4] on a generalization of Shannon information measures to arbitrary $\sigma$-fields) or if $H(K_{t})<\infty$ (this follows by approximating $K_{t}$ by finite-alphabet random variables). Now suppose that process $(K_{t})_{t\in\mathbb{Z}}$ is stationary and $H(K_{t})<\infty$. We have $h=H(K_{t+1}|K_{-\infty}^{t})$ by (82) and $v=0$ by Proposition 9. Since $H(K_{k+1}^{k+s}|K_{1}^{t})\geq H(K_{k+1}^{k+s}|K_{-\infty}^{k})=hs$ for $k\geq t$, we may write
$\sup_{k\geq t}\frac{H(X_{k+1}^{k+s}|X_{1}^{t})}{s}$ $\geq\inf_{k\geq t}\frac{H(K_{k+1}^{k+s}|K_{1}^{t})}{s}+C_{7}\sup_{k\geq t}\frac{V(\mathcal{K}_{k+1}^{k+s}\setminus\mathcal{K}_{1}^{t})}{s}$
$\geq h+C_{7}\sup_{k\geq t}\frac{V(\mathcal{K}_{k+1}^{k+s}\setminus\mathcal{K}_{1}^{t})}{s}.$ (83)
Further, by Proposition 9, we may lower bound
$\sup_{k\geq t}\frac{V(\mathcal{K}_{k+1}^{k+s}\setminus\mathcal{K}_{1}^{t})}{s}\geq\frac{V(t+s)-V(t)}{s}\geq\Delta V(t+s-1)\geq\Delta V(t+s).$ (84)
Hence we obtain claim (77). ∎
3.3 Hilberg‘s law implies neural scaling
Finally, we will show that a differential Hilberg law implies the neural scaling law for an arbitrary stochastic process with a finite entropy rate.
Proposition 15.
Let $(X_{t})_{t\in\mathbb{Z}}$ be an arbitrary stochastic process such that $h<\infty$ and we have a version of Hilberg‘s law of form
$\sup_{k\geq t}\frac{H(X_{k+1}^{k+s}|X_{1}^{t})}{s}-h$ $\geq C_{9}(t+s)^{\beta-1}.$ (85)
for a certain $\beta\in(0,1)$. Let $Q_{tnc}$ be a random probability measure that satisfies
$H(Q_{tnc}|X_{1}^{t})$ $\leq C_{9}c,$ (86)
$H(Q_{tnc})$ $\leq C_{9}n.$ (87)
Define also
$s_{\max}(t,n,c)$ $:=\min\left\{\frac{t\sqrt{\frac{ct^{-\beta}}{1-\beta}}}{1-\sqrt{\frac{ct^{-\beta}}{1-\beta}}},\left(\frac{n}{1-\beta}\right)^{\frac{1}{\beta}}\right\}.$ (88)
Then for $s\leq s_{\max}(t,n,c)$ we have the neural scaling law of form
$\sup_{k\geq t}\frac{\operatorname{\mathbf{E}}(-\log Q_{tnc}(X_{k+1}^{k+s}))}{s}-h$
$\geq C_{9}\max\left\{t^{\beta-1}\left(\frac{1-\sqrt{\frac{ct^{-\beta}}{1-\beta}}}{1+\sqrt{\frac{ct^{-\beta}}{1-\beta}}}\right)^{1-\beta}-\frac{c}{t}\left(\frac{1-\sqrt{\frac{ct^{-\beta}}{1-\beta}}}{2\sqrt{\frac{ct^{-\beta}}{1-\beta}}}\right),\frac{2^{\beta}-1+\beta}{2}\left(\frac{n}{1-\beta}\right)^{1-\frac{1}{\beta}}\right\}.$ (89)
Proof.
Fix a $\beta\in(0,1)$. We denote the function
$f(s)=f(s,t,c):=(t+s)^{\beta-1}-\frac{c}{s}$ (90)
for $t,c\geq 0$. Let $r=r(t,c):=\operatorname*{arg\,max}_{s>0}f(s,t,c)$. We have
$0=\left.\frac{df(s)}{ds}\right|_{s=r}=-(1-\beta)(t+r)^{\beta-2}+\frac{c}{r^{2}}.$ (91)
Consequently, $(1-\beta)(t+r)^{\beta-2}r^{2}=c$.
Assume first that $t>0$. Then we may bound
$(1-\beta)t^{\beta-2}r^{2}\geq(1-\beta)(t+r)^{\beta-2}r^{2}\geq(1-\beta)t^{\beta}\left(\frac{r}{t+r}\right)^{2}.$ (92)
Hence we obtain the sandwich bound $r_{0}\leq r\leq r_{2}$, where
$r_{0}=r_{0}(t,c)$ $:=ty,$ (93)
$r_{2}=r_{2}(t,c)$ $:=\frac{ty}{1-y},$ $y$ $:=\sqrt{\frac{ct^{-\beta}}{1-\beta}}.$ (94)
Function $f(s)$ is increasing for $s\leq r$ and decreasing for $s\geq r$. Suppose that $s\leq r_{2}$. For $q=\left\lceil r_{2}/s\right\rceil$, we may bound
$r\leq sq\leq s\left(\frac{r_{2}}{s}+1\right)\leq r_{2}+s\leq 2r_{2}$ (95)
so we can also bound
$f(sq,t,c)$ $\geq f(2r_{2})=t^{\beta-1}\left(\frac{1-y}{1+y}\right)^{1-\beta}-\frac{c}{t}\left(\frac{1-y}{2y}\right).$ (96)
Now assume that $t=0$ and $c=n$. Then
$r=r(0,n):=\left(\frac{n}{1-\beta}\right)^{\frac{1}{\beta}}$ (97)
Suppose that $s\leq r$. For $q=\left\lceil r/s\right\rceil$, we may bound
$r\leq sq\leq s\left(\frac{r}{s}+1\right)\leq r+s\leq 2r$ (98)
so we can also bound
$f(sq,0,n)$ $\geq f(2r)=\frac{2^{\beta}-1+\beta}{2}\left(\frac{n}{1-\beta}\right)^{1-\frac{1}{\beta}}.$ (99)
This completes the analysis of function $f(s)$ that will be needed next.
Now we proceed to the main part of the proof. By an iterated application of inequality (21), for any $q\in\mathbb{N}$, we have inequality
$\sup_{k\geq t}\frac{H(X_{k+1}^{k+s}|Q_{tnc})}{s}\geq\sup_{k\geq t}\frac{H(X_{k+1}^{k+sq}|Q_{tnc})}{sq}.$ (100)
By contrast, by inequality (25), conditions (86) and (87) imply inequality
$H(X_{k+1}^{k+s}|Q_{tnc})\geq\max\left\{H(X_{k+1}^{k+s}|X_{1}^{t})-C_{9}c,H(X_{k+1}^{k+s})-C_{9}n\right\}.$ (101)
Consequently, for $s\leq r_{2}(t,c)$ and $q=\left\lceil r_{2}(t,c)/s\right\rceil$, condition (85) yields
$\sup_{k\geq t}\frac{H(X_{k+1}^{k+s}|Q_{tnc})}{s}-h$ $\geq\sup_{k\geq t}\frac{H(X_{k+1}^{k+sq}|Q_{tnc})}{sq}-h\geq\sup_{k\geq t}\frac{H(X_{k+1}^{k+sq}|X_{1}^{t})-C_{9}c}{sq}-h$
$\geq C_{9}f(sq,t,c)\geq C_{9}\left(t^{\beta-1}\left(\frac{1-y}{1+y}\right)^{1-\beta}-\frac{c}{t}\left(\frac{1-y}{2y}\right)\right).$ (102)
Complementing the above inequality with the analogous development for conditions $s\leq r(0,n)$ and $q=\left\lceil r(0,n)/s\right\rceil$ yields
$\sup_{k\geq t}\frac{H(X_{k+1}^{k+s}|Q_{tnc})}{s}-h$ $\geq\sup_{k\geq t}\frac{H(X_{k+1}^{k+sq}|Q_{tnc})}{sq}-h\geq\sup_{k\geq t}\frac{H(X_{k+1}^{k+sq})-C_{9}n}{sq}-h$
$\geq C_{9}f(sq,0,n)\geq C_{9}\frac{2^{\beta}-1+\beta}{2}\left(\frac{n}{1-\beta}\right)^{1-\frac{1}{\beta}}.$ (103)
Hence, by the source coding inequality (29), we recover the claim (89). ∎
4 Conclusion
Formalizing prior results in machine learning [50, 61, 64, 14, 70, 71] and extending earlier results in information theory and quantitative linguistics [23, 25, 26, 28], we have derived of a chain of deductive implications that connect Zipf‘s law to neural scaling. By isolating the discrete steps that go through Heaps‘ law and Hilberg‘s hypothesis and by giving explicit assumptions needed for each step, we have attempted to clarify which aspects of natural language properties are responsible for the power-law behavior observed in foundation models. Our results show that once vocabulary growth exhibits a power-law growth and once block entropy inherits this scaling, then the constraints imposed by limited data, parameters, and compute naturally produce a generic neural scaling law.
Our theoretical derivation of underparameterization as the optimal regime contrasts sharply with the empirical success of overparameterized models [51]. This discrepancy suggests that naive parameter counts may overestimate the effective capacity of real-valued neural networks, as larger models may use fewer digits of the binary expansion per real-valued parameter. If this holds true then the received concept of overparameterization may be misleading and determining the true number of degrees of freedom of a foundation model may not be so straightforward. Likewise, our purely information-theoretic modeling of compute—via entropy constraints—offers only a coarse abstraction of training dynamics. We suppose that a refined treatment may require tools from resource-bounded algorithmic information theory.
Finally, the reduction of the neural scaling law to Hilberg‘s hypothesis shifts attention to the deeper question of why power-law scaling of entropy arises in natural language at all. One simple possibility is that Hilberg‘s law emerges for Zipfian vocabularies with word shapes that are sufficiently arbitrary or algorithmically random, as suggested by the Santa Fe decomposition (7). Establishing such a connection rigorously requires a more detailed analysis of universal coding, extending those of works [25, 26, 27]. We hope that the framework developed here provides a baseline for such investigations and that future work will refine the idealizations that we have identified.
AI assistance
Several paragraphs in the introduction and in the conclusion have been co-drafted by the author and ChatGPT and critically post-edited by the author. No reference was suggested by the AI assistant.
Funding
This work received no external funding.

### v2

@@
 Key words: neural scaling law, Zipf‘s law, Heaps‘ law, Hilberg‘s hypothesis, Santa Fe processes
-MSC 2020: 94A17, 60G10
+MSC 2020: 94A17, 60G10, 91F20, 68T07
 1 Introduction
@@
 This emerging perspective suggests that some remarkable empirical regularities in large-scale deep learning need not be explained solely by architectural, optimization, or hardware considerations, but instead reflect intrinsic statistical constraints of natural texts—human linguistic data.
-In the present paper, we aim to provide an actually simple but systematic derivation of this chain of implications that extends earlier results in information theory and quantitative linguistics [23, 25, 26, 28].
+In the present paper, we aim to provide an actually simple but systematic derivation of this chain of implications that extends earlier results in information theory, probability theory, and quantitative linguistics.
 We will prove formal theorems about as general stochastic processes as possible rather than experiment with particular empirical data.
 Our goal is to consolidate knowledge of several scientific disciplines by means of mathematical deduction.
+We try to avoid a too abstract formalism to make this paper more accessible.
 The conceptual trajectory is as follows.
 Beginning with Zipf‘s law for word frequencies, we derive Heaps‘ law for vocabulary growth.
-From Heaps‘ law we extract a formal version of Hilberg‘s hypothesis on the sublinear growth of block entropy.
-Finally, we show that Hilberg‘s hypothesis leads to the neural scaling that ties model performance to training data size, model size, and amount of training compute.
-In parallel, we will discuss an example of Santa Fe processes—toy stochastic sources that exhibit these statistical laws.
-Each link in this chain introduces its own assumptions and modeling choices.
-By isolating these steps, we hope to illuminate where the derivations rely on idealizations and how they might be strengthened.
+From Heaps‘ law we extract Hilberg‘s hypothesis on the sublinear growth of block entropy.
+Finally, we show that Hilberg‘s hypothesis leads to the neural scaling that ties model performance to the amounts of training data, model parameters, and training compute.
+In parallel, we discuss Santa Fe processes—toy stochastic sources that exhibit these statistical laws.
+Each link in this chain introduces its own assumptions.
+By isolating these steps, we hope to illuminate where the derivations might be strengthened in the future.
 Notations.
@@
 We adopt that $X_{1}^{0}$ is the empty string and $\mathcal{X}_{1}^{0}$ is the empty set.
-Notation $H(X):=\operatorname{\mathbf{E}}(-\log P(X))$ stands for the Shannon entropy of a random variable and $V(\mathcal{X}):=\operatorname{\mathbf{E}}\operatorname{\#}\mathcal{X}$ denotes the expected cardinality of a random set.
+Expectation with respect to probability measure $P$ is denoted $\operatorname{\mathbf{E}}X:=\int XdP$, $H(X):=\operatorname{\mathbf{E}}(-\log P(X))$ is the Shannon entropy, and $V(\mathcal{X}):=\operatorname{\mathbf{E}}\operatorname{\#}\mathcal{X}$ is the expected cardinality.
 The conditional entropy is H(X|Q):=H(X,Q)H(Q)H(X|Q):=H(X,Q)-H(Q), the mutual information is I(X;Y):=H(X)H(X|Y)I(X;Y):=H(X)-H(X|Y), and the conditional mutual information is I(X;Y|Q):=H(X|Q)H(X|Y,Q)I(X;Y|Q):=H(X|Q)-H(X|Y,Q).
-We also use notations
+We also use
 $f(x)\stackrel{{\scriptstyle*}}{{<}}g(x)$ $\iff f(x)\leq cg(x)\text{ for a $c>0$},$ (1)
@@
 The oldest known of quantitative linguistic laws, Zipf‘s law asserts that, for texts in natural language, the frequency of the $k$-th most common type of word decays approximately as $k^{-\alpha}$, where $\alpha\approx 1$.
-This regularity was noticed over one hundred years ago [31, 17, 90].
+This regularity was noticed over one century ago [33, 19, 98].
 An empirical study of this law across one hundred languages can be found in [67].
@@
 Historical references can be found in [63].
-The departure point of a statistically informed theory is [52, 4].
-In spite of sheer literature coverage, there is no single explanation of Zipf‘s law.
-The law can be explained by diverse mechanisms ranging from monkey-typing [62, 68], through preferential attachment [80] and game theory [41], to potential links with semantics and information theory [26].
-What matters for our purposes is that stochastic processes $(K_{t})_{t\in\mathbb{Z}}$ over natural numbers with the Zipf probability distribution
+The departure point of a statistically informed theory of Zipf‘s law is formed by references [54, 55, 78, 5].
+In spite of sheer literature coverage (or maybe because of that), there is no single explanation of Zipf‘s law.
+The law can be explained by diverse mechanisms ranging from monkey-typing [66, 72], through preferential attachment [87], also known as the Chinese restaurant process [3, page 92], to potential links with game theory [43], semantics, and information theory [28].
+What matters for our purposes is that stochastic processes $(K_{t})_{t\in\mathbb{Z}}$ over natural numbers with the approximate Zipf probability distribution
 $P(K_{t}=k)\stackrel{{\scriptstyle*}}{{=}}k^{-\alpha},\quad\alpha>1,$ (4)
-provide a compact description of heavy-tailed data distributions. Observation (4) will be our departure point.
+provide a compact description of heavy-tailed data distributions. Statistical law (4) will be our departure point.
 Heaps‘ law.
@@
 $V(\mathcal{X}_{1}^{t})\stackrel{{\scriptstyle*}}{{=}}t^{\beta},\quad 0<\beta<1.$ (5)
-Heaps‘ law is widely viewed as a direct consequence of Zipf‘s law with $\beta=1/\alpha$ for $X_{t}=K_{t}$ [62] but the linguistic phenomenology and the statistical theory are more complicated [52, 4, 66, 67, 39, 35, 20, 29].
-Understanding precisely what form Zipf‘s law implies a particular form of Heaps‘ law is an important step in our approach.
+Heaps‘ law is widely viewed as a direct consequence of Zipf‘s law with $\beta=1/\alpha$ for $X_{t}=K_{t}$ [66] but the linguistic phenomenology and the statistical theory are more complicated [54, 55, 5, 70, 71, 41, 37, 22, 16, 31].
+Heaps‘ law has been also studied for large language models [90, 58].
+Understanding precisely what form of Zipf‘s law implies a particular form of Heaps‘ law is an important step in our approach.
 Hilberg‘s hypothesis.
@@
 Thus, we may interpret that sequence $(Z_{k})_{k\in\mathbb{N}}$ expresses some unbounded immutable general knowledge which is only partially accessed and described in finite texts.
-[…]
+Contrary to intuitions about complex structures, Hilberg‘s law arises also in highly simplified settings of Santa Fe processes.
+In particular, it suffices to assume that narration $(K_{t})_{t\in\mathbb{Z}}$ is an IID source with the Zipf distribution (4) and knowledge $(Z_{k})_{k\in\mathbb{N}}$ is a sequence of independent fair coin flips, independent of narration $(K_{t})_{t\in\mathbb{Z}}$.
+Under these conditions, we obtain Heaps‘ law (5) and Hilberg‘s law (6) with $\beta=1/\alpha$ [27, 28].
 Neural scaling.
@@
 A particularly salient empirical finding is that the predictive performance of these models improves with scale in a power-law fashion.
-The neural scaling law characterizes how the loss function of a foundation model—typically measured by the cross entropy rate on a test dataset—decreases as training data $t$, model size $n$, and compute $c$ increase [51, 44, 46, 48, 58].
+The neural scaling law characterizes how the loss function of a foundation model—typically measured by the cross entropy rate on a test dataset—decreases as training data $t$, model size $n$, and compute $c$ increase [53, 46, 48, 50, 76, 62].
 Simplifying particular empirical observations and ignoring complex interactions between $t$, $n$, and $c$, this power-law relationship can be approximated as
@@
 for a fixed $s<\infty$, where
-$h(s,t,n,c):=\frac{\operatorname{\mathbf{E}}(-\log Q_{tnc}(X_{t+1}^{t+s}))}{s}$ (11)
-is the expected cross entropy rate of a foundation model $Q_{tnc}$ tested on data $X_{t+1}^{t+s}$ and trained on data $X_{1}^{t}$ with the amount of parameters $n$ and the amount of compute $c$.
+$h(s,t,n,c):=\sup_{k\geq t}\frac{\operatorname{\mathbf{E}}(-\log Q_{tnc}(X_{k+1}^{k+s}))}{s}$ (11)
+is the worst-case expected cross entropy rate of a foundation model $Q_{tnc}$ tested on data $X_{k+1}^{k+s}$ and trained on data $X_{1}^{t}$ with the amount of parameters $n$ and the amount of compute $c$.
 The exponents satisfy $\gamma_{i}>0$.
-Empirically, it is confirmed that $\gamma_{T}>\gamma_{N}$ [51].
+Empirically, it is confirmed that $\gamma_{T}>\gamma_{N}$ [53, 50, 76].
 This case is called the overparameterization.
@@
 We will systematically prove connections among four statistical laws in particular settings that have not been discussed so far.
-The present paper develops and strengthens the ideas from an earlier and unpublished attempt to attack this topic [28].
+The present paper develops and strengthens the ideas from an earlier unpublished attempt to attack this topic [30].
 We would like to supply a simple-minded baseline that, in contrast to [30], takes into account also the effect of limited amount of training compute $c$.
 The main contributions of the present paper are organized as a chain of following derivations:
-Zipf‘s law $\stackrel{{\scriptstyle(A)}}{{\Rightarrow}}$ Heaps‘ law $\stackrel{{\scriptstyle(B)}}{{\Rightarrow}}$ Hilberg‘s law $\stackrel{{\scriptstyle(C)}}{{\Rightarrow}}$ neural scaling.
-Each of the above implications involves different assumptions but we strive at the most general settings such as non-stationary processes.
-We remark that the conditions for implication $(C)$ (non-stationarity) are more general than that for result $(B)$ (stationarity and Santa Fe decomposition), and those are more general than the requirements for implication $(A)$ (stationarity and mixing).
-We also notice that the derivation of the neural scaling law rests on the differential Heaps and Hilberg laws (15)–(16), which strengthen the plain Heaps and Hilberg laws (5)–(6).
+Zipf‘s law $\stackrel{{\scriptstyle(A)}}{{\implies}}$ Heaps‘ law $\stackrel{{\scriptstyle(B)}}{{\implies}}$ Hilberg‘s law $\stackrel{{\scriptstyle(C)}}{{\implies}}$ neural scaling.
+Each of the above implications involves different assumptions but we strive at the most general settings—such as non-stationary processes.
+In particular, the conditions for implication $(C)$ (arbitrary non-stationarity) are more general than that for result $(B)$ (stationarity and Santa Fe decomposition), and those are more general than the requirements for implication $(A)$ (stationarity and mixing).
+We observe that the derivation of neural scaling rests on the differential Heaps and Hilberg laws (15)–(16), which strengthen the plain Heaps and Hilberg laws (5)–(6).
 Noticing a practical need for such differential laws seems a novel idea of this paper.
+It would be nice to combine this idea with the powerful IID framework by Karlin [54] in the future, possibly also extending it from IID to mixing or ergodic processes.
 Let us begin with result $(C)$, to be followed by $(B)$ and $(A)$. Let $Q_{tnc}$ be a random probability measure that satisfies
@@
 (The validity of these assumptions is discussed in the next paragraph.) In the result $(C)$, for a sufficiently small test sample length $s$, we derive the neural scaling law of form
-$\sup_{k\geq t}\frac{\operatorname{\mathbf{E}}(-\log Q_{tnc}(X_{k+1}^{k+s}))}{s}-h$
+$h(s,t,n,c)-h$
 $\stackrel{{\scriptstyle*}}{{>}}\max\left\{t^{\beta-1}\left(\frac{1-\sqrt{\frac{ct^{-\beta}}{1-\beta}}}{1+\sqrt{\frac{ct^{-\beta}}{1-\beta}}}\right)^{1-\beta}-\frac{c}{t}\left(\frac{1-\sqrt{\frac{ct^{-\beta}}{1-\beta}}}{2\sqrt{\frac{ct^{-\beta}}{1-\beta}}}\right),\frac{2^{\beta}-1+\beta}{2}\left(\frac{n}{1-\beta}\right)^{1-\frac{1}{\beta}}\right\}.$ (14)
@@
 $\sup_{k\geq t}\frac{V(\mathcal{K}_{k+1}^{k+s}\setminus\mathcal{K}_{1}^{t})}{s}$ $\stackrel{{\scriptstyle*}}{{>}}(t+s)^{\beta-1}.$ (16)
-Finally, we show that this form Heaps‘ law is satisfied by sufficiently strongly mixing narrations with sufficiently heavy-tailed marginal distributions.
-In particular, we derive result $(A)$ for a stationary process $(K_{t})_{t\in\mathbb{Z}}$ with a mixing condition $\Psi<\infty$, where
-$\Psi$ $:=\sup_{t,k}\frac{\sum_{i=1}^{t}\sum_{j=1}^{t}P(K_{i}=k,K_{j}=k)-tP(K_{0}=k)}{t(t-1)(P(K_{0}=k))^{2}},$ (17)
-and an approximate Zipf law
+Finally, we show that this form of Heaps‘ law is satisfied by sufficiently strongly mixing narrations with sufficiently heavy-tailed marginal distributions.
+In particular, we derive result $(A)$ for a stationary process $(K_{t})_{t\in\mathbb{Z}}$ such that
+$\sum_{k\in\mathbb{N}}P(K_{0}=k){\bf 1}{\left\{P(K_{0}=k)\leq p\right\}}$ $\stackrel{{\scriptstyle*}}{{>}}p^{1-\beta},$ (17)
+$\frac{P(K_{t}=k|K_{0}=k)}{P(K_{0}=k)}$ $\stackrel{{\scriptstyle*}}{{<}}1.$ (18)
+In particular, condition (17) holds for an approximate Zipf law
 $P(K_{0}=k)$ $\stackrel{{\scriptstyle*}}{{=}}k^{-1/\beta}.$ (19)
+Condition (18) is satisfied for IID and finite-state processes. It is implied by condition $\psi^{*}(1)<\infty$ in the terminology of Bradley [13]. We also note that
+$\lim_{t\to\infty}\frac{P(K_{t}=k|K_{0}=k)}{P(K_{0}=k)}=1$ (20)
+holds for any stationary strongly mixing process $(K_{t})_{t\in\mathbb{Z}}$. For a comprehensive survey of various mixing conditions, we refer to [13].
 Open problems.
 Our analysis highlights some gaps that need to be addressed in the future research:
+  • Tightness of the derived bounds. Suppose that the neural scaling law of form (14) holds. Then we derive inequalities
+    $\gamma_{T}$ $\leq 1-\beta,$ (21)
+    $\gamma_{N}$ $\leq\frac{1}{\beta}-1.$ (22)
+    In particular, for the PPM-based estimate $\beta\approx 0.8$ reported by Takahira et al.
+[88], we obtain $\gamma_{T}\leq 0.2$ and $\gamma_{N}\leq 0.25$.
+This is quite a loose upper bound of the values $\gamma_{T}\approx 0.095$ and $\gamma_{N}\approx 0.076$ reported by Kaplan et al.
+[53].
+We note that these estimates were obtained for corpora of a different magnitude—gigabytes of tokens in the case of [88] to be contrasted with terabytes of tokens in the case of [53].
+Further data analysis is needed to explain the large gap between the empirical values of exponents $\gamma_{T}$ and $\gamma_{N}$ and their upper bounds stemming from known estimates of the Hilberg exponent $\beta$.
+It is plausible that $\beta$ is substantially larger than $0.8$ for internet-sized corpora.
   • Modeling bounds on compute.
@@
   • Overparameterization.
-The result stated in equation (14) suggests theoretical optimality of underparameterization $\gamma_{T}\leq\gamma_{N}$, contrary to the empirical optimality of overparameterization $\gamma_{T}>\gamma_{N}$ [51].
-Indeed, suppose that (14) holds.
-Then we derive
-    $\gamma_{T}$ $\leq 1-\beta,$ (19)
-    $\gamma_{N}$ $\leq\frac{1}{\beta}-1.$ (20)
-    Thus understanding why overparameterized models perform empirically better requires an explanation.
+The derived inequalities (21)–(22) suggest theoretical optimality of underparameterization $\gamma_{T}\leq\gamma_{N}$, contrary to the empirical optimality of overparameterization $\gamma_{T}>\gamma_{N}$ [53, 50, 76].
+Thus understanding why overparameterized models perform empirically better may require an explanation.
 One such explanation may be that condition (13) does not capture the fact that a single real-valued parameter (a single weight) may carry potentially an infinite amount of information (theoretically, it is an infinite sequence of bits).
@@
 The internal randomness of these tokens can be potentially connected to semantic or linguistic considerations [27, 28, 68].
-In particular, we may expect that Hilberg‘s law arises if we have Zipf‘s law for words that have sufficiently arbitrary shapes, cf.
-[22].
-Proving such a claim formally requires a longer excursion to universal coding and involves results that exceed those of papers [25, 26, 27].
+In particular, we may expect that Hilberg‘s law is equivalent to Zipf‘s law for words that have sufficiently arbitrary shapes.
+Arbitrariness of word shapes is one of the classical tenets of linguistics [24].
+Proving this equivalence formally requires a longer excursion to universal coding and involves results that exceed those of papers [27, 28, 29].
 For this reason, we postpone this theoretical development to another article.
@@
 Section 2.3 handles the block entropy and the expected block cardinality for stationary processes.
-Section 2.4 analyzes the vocabulary spectrum elements for IID processes.
+It also introduces spectrum elements and bounds the rate of hapaxes.
+Section 2.4 analyzes the spectrum elements for IID processes.
 Section 3 develops the main chain of implications.
@@
 For the textbook treatment and the background, we refer to [20, 96].
-A handy tool is also a generalization of Shannon information measures to arbitrary $\sigma$-fields [26, Sections 5.3 and 5.4].
+A handy tool is also a generalization of Shannon information measures to arbitrary $\sigma$-fields [28, Sections 5.3 and 5.4].
 In general, both the Shannon entropy and the expected cardinality are subadditive and enjoy the triangle inequality.
@@
 $\geq\operatorname{\mathbf{E}}\operatorname{\mathbf{E}}\left(-\log P(X|Q)\middle|Q\right)$
-$=\operatorname{\mathbf{E}}\left(-\log P(X|Q)\right)$
-$=H(X|Q)$
-$\geq H(X)-H(Q),$ (30)
+$=\operatorname{\mathbf{E}}\left(-\log P(X|Q)\right)=H(X|Q),$ (32)
 where we have used the law of total expectation and non-negativity of the Kullback-Leibler divergence. ∎
@@
 Proof.
-See [26, §5.2]. In general, we have
+See [28, §5.2]. In general, we have
 $H(t)$ $=H(X_{1}^{t})\geq 0,$ (43)
@@
 ∎
-Besides the expected number of all types $V(t)$, let us introduce expected spectrum elements $V(t|m)$, defined as the expected number of types that occur exactly $m$ times.
+Besides the expected number of all types $V(t)$, let us introduce spectrum elements $V(t|m)$, defined as the expected number of types that occur exactly $m$ times [54, 55, 5].
 Definition 3.
-The expected spectrum elements are defined as
-$V(t|m):=\sum_{k}P\left(\sum_{i=1}^{t}{\bf 1}{\left\{K_{i}=k\right\}}=m\right).$ (51)
+The spectrum elements are defined as
+$V(t|m)$ $:=\sum_{k\in\mathbb{N}}P\left(\sum_{i=1}^{t}{\bf 1}{\left\{K_{i}=k\right\}}=m\right),$ $1$ $\leq m\leq t.$ (53)
 In particular, $V(t|1)$ is the expected number of types that occur once.
@@
 So far, the statements for Shannon entropy and expected cardinality were mirror-like.
-However, for more specific processes, the analogy between these two functionals is rather as follows: The expected cardinality applies to IID processes as the Shannon entropy applies to exchangeable Santa Fe processes.
-This analogy will generalized formally in Section 3.2.
-Now let us comment on the expected cardinality for IID processes, as it is simpler to analyze.
-We notice that the expected cardinality for IID sources enjoys further properties: It is a Hausdorff sequence—a discrete-time analog of the Bernstein function.
-Combining creatively several sources such as [4, 2, 76], the development is as follows. By the Bernoulli scheme, for an IID process $(K_{t})_{t\in\mathbb{Z}}$, we have
-$V(t)$ $=\sum_{k}(1-(1-p_{k})^{t}),$ (54)
-$V(t|m)$ $=\begin{cases}\sum_{k}\binom{t}{m}p_{k}^{m}(1-p_{k})^{t-m},&1\leq m\leq t,\\ 0,&m>t,\end{cases}$ (55)
-where $p_{k}:=P(K_{t}=k)$. We notice consistency conditions
-$\sum_{m=1}^{\infty}V(t|m)$ $=V(t),$ $\sum_{m=1}^{\infty}mV(t|m)$ $=t.$ (56)
-Moreover, since $\Delta(1-p_{k})^{t}=-p_{k}(1-p_{k})^{t}$, we also observe that the spectrum elements can be computed from the expected number of types by taking differences,
-$V(t|m)$ $=\begin{cases}(-1)^{m+1}\binom{t}{m}\Delta^{m}V(t-m),&1\leq m\leq t,\\ 0,&m>t.\end{cases}$ (57)
-Expression (57) reminds of a known concept.
-Definition 4 (cf. [2]).
+However, for more specific processes, the analogy between these two functionals is rather as follows: The expected cardinality applies to IID processes in a similar fashion as the Shannon entropy applies to exchangeable Santa Fe processes.
+This analogy will be generalized in Section 3.2.
+Now we consider the expected cardinality for IID processes, being simpler to analyze.
+We notice that the expected cardinality for IID sources enjoys further properties: It is a Hausdorff sequence—a discrete-time analog of a Bernstein function.
+This observation nicely complements the results by Karlin [54] for IID and Poisson cases.
+The development is as follows.
+Definition 4.
 A sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ is called a Hausdorff sequence if $v(t)\geq 0$ and $(-1)^{m+1}\Delta^{m}v(t)\geq 0$ for all $m\in\mathbb{N}$ and $n\in\mathbb{N}\cup\left\{0\right\}$.
-A sequence $u:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ is called completely alternating if $u(t)=\Delta v(t)$ for a certain Hausdorff sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$.
-We call a Hausdorff sequence $v(t)$ standard if $v(0)=0$ and $\Delta v(0)=1$.
-The name Hausdorff sequence is non-standard itself.
+A sequence $u:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ is called completely alternating if $u(t)=\Delta v(t)$ for a certain Hausdorff sequence $v(t)$.
+We call a sequence $v(t)$ standard if $v(0)=0$ and $\Delta v(0)=1$.
+Remark: The name Hausdorff sequence is non-standard itself.
 We have coined it by an analogy to the standard term Bernstein function, which is the continuous time-analog, applying derivatives rather than differences [83].
@@
 It is a discrete-time analog of the Bernstein theorem on completely monotone functions [11], also known as the Lévy-Khintchine representation in probability [61].
-Proposition 11 ([2]).
-A sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ is a Hausdorff sequence if and only if there exists a unique non-negative measure $F$ on $[0,1]$ such that
-$v(t)$ $=tF(\left\{0\right\})+\int_{(0,1)}\frac{1-(1-p)^{t}}{p}dF(p)+F(\left\{1\right\}),$ (58)
-$\Delta v(t)$ $=F(\left\{0\right\})+\int_{(0,1)}(1-p)^{t}dF(p).$ (59)
-Additionally, we see that a Hausdorff sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ is standard if and only if $F(\left\{1\right\})=0$ and the measure $F$ is a probability measure, namely,
-$F(\left\{0\right\})+\int_{(0,1)}dF(p)=1.$ (60)
-We call measure $F$ the Hausdorff measure.
-We notice that the Newton formula $(1-\Delta)^{r}=\sum_{m=0}^{r}\binom{r}{m}(-\Delta)^{m}$ can be rewritten in the following form, which resembles the Taylor expansion,
-$v(t-r)$ $=\sum_{m=0}^{r}(-1)^{m}\binom{r}{m}\Delta^{m}v(t-m).$ (61)
-By formula (61), consistency conditions analogous to (56) can be verified easily. Namely, a Hausdorff sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ with spectrum elements
-$v(t|m)$ $:=\begin{cases}(-1)^{m+1}\binom{t}{m}\Delta^{m}v(t-m),&1\leq m\leq t,\\ 0,&m>t.\end{cases}$ (62)
-satisfies conditions $\sum_{m=1}^{t}v(t|m)=v(t)$ and $\sum_{m=1}^{t}mv(t|m)=t$ if and only if it is standard.
-To complete this discussion, the expected number of types $V(t)$ is a standard Hausdorff sequence and it has a Hausdorff probability measure supported on discrete points, namely,
-$F(A)=\sum_{k}p_{k}{\bf 1}{\left\{p_{k}\in A\right\}}.$ (63)
+Proposition 11.
+A sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ is a Hausdorff sequence if and only if there exists a unique non-negative measure $\tilde{v}$ on $[0,1]$ such that
+$v(t)$ $=t\tilde{v}(\left\{0\right\})+\int_{(0,1)}\frac{1-(1-p)^{t}}{p}d\tilde{v}(p)+\tilde{v}(\left\{1\right\}),$ (56)
+$\Delta v(t)$ $=\tilde{v}(\left\{0\right\})+\int_{(0,1)}(1-p)^{t}d\tilde{v}(p).$ (57)
+Remark: We call measure $\tilde{v}$ the Hausdorff measure of sequence $v(t)$.
+A Hausdorff sequence $v(t)$ is standard if and only if $\tilde{v}(\left\{1\right\})=0$ and the measure $\tilde{v}$ is a probability measure.
+Proof.
+See [44, 2]. ∎
+Definition 5.
+For an arbitrary sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$, we define its Taylor elements
+$v(t\|m)$ $:=(-1)^{m+1}\binom{t}{m}\Delta^{m}v(t-m),$ $1$ $\leq m\leq t.$ (58)
+We notice that $v(t\|m)\geq 0$ if and only if sequence $v(t)$ is a Hausdorff sequence.
+Proposition 12.
+The Taylor elements satisfy consistency conditions
+$\sum_{m=1}^{\infty}v(t\|m)$ $=v(t),$ $\sum_{m=1}^{\infty}mv(t\|m)$ $=t.$ (59)
+if and only if sequence $v(t)$ is standard.
+Proof.
+The claim follows by the Newton formula $(1-\Delta)^{r}=\sum_{m=0}^{r}\binom{r}{m}(-\Delta)^{m}$, written as
+$v(t-r)$ $=\sum_{m=0}^{r}(-1)^{m}\binom{r}{m}\Delta^{m}v(t-m),$ (60)
+resembling the Taylor expansion. It suffices to consider (60) for $r=t$ and r=t1r=t-1. ∎
+Now, without loss of generality, let us assume that the alphabet of an IID process $(K_{t})_{t\in\mathbb{Z}}$ is the set of natural numbers.
+By the Bernoulli scheme, the expected number of types and the spectrum elements defined via (53) equal
+$V(t)$ $=\sum_{k\in\mathbb{N}}(1-(1-p_{k})^{t}),$ (61)
+$V(t|m)$ $=\sum_{k\in\mathbb{N}}\binom{t}{m}p_{k}^{m}(1-p_{k})^{t-m},$ $1$ $\leq m\leq t,$ (62)
+where $p_{k}:=P(K_{0}=k)$. By formula (62) and identity $\Delta(1-p_{k})^{t}=-p_{k}(1-p_{k})^{t}$, the spectrum elements $V(t|m)$ and the Taylor elements of $V(t)$ are equal,
+$V(t|m)=V(t\|m).$ (63)
+Consequently, sequence $V(t)$ is a Hausdorff sequence.
+Moreover, sequence $V(t)$ is standard and spectrum elements $V(t|m)$ satisfy consistency conditions analogous to (59).
+The Hausdorff measure of $V(t)$ is an atomic probability measure and assumes form
+$\tilde{V}(A)=\sum_{k\in\mathbb{N}}p_{k}{\bf 1}{\left\{p_{k}\in A\right\}}.$ (64)
 In Section 3.1, we approximate this measure by a non-atomic measure so as to derive two symmetric bounds for the conditional Heaps law from a bound for Zipf‘s law.
@@
 3.1 Zipf‘s law implies Heaps‘ law
-Let us recall from Section 2.4 that the expected number of hapaxes for IID processes satisfies
-$V(t+1|1)=(t+1)\Delta V(t).$ (64)
-The growth rate of the number of hapaxes $V(t+1|1)$ is controlled by the tail of the marginal distribution.
-In fact, there are two simple symmetric conditions on the tail of the marginal distribution that sandwich bound the number of hapaxes and lead to a differential Heaps law.
-We continue notation $p_{k}:=P(K_{t}=k)$ in the following.
+By formulas (58) and (63), the expected rate of hapaxes for IID processes equals
+$\frac{V(t|1)}{t}=\Delta V(t-1),$ (65)
+which can be contrasted with the more general inequality (54) for stationary sources.
+Thus some bounds for difference $\Delta V(t)$ can be obtained by showing that the rate of hapaxes $V(t|1)/t$ is controlled by the tail of the marginal distribution.
+In fact, there are two symmetric conditions on the tail of the marginal distribution that sandwich the number of hapaxes and lead to a differential Heaps law.
+Similar bounds, though asymptotic and covering only the IID case, were discussed by Karlin [54].
+In the following, we use $p_{k}:=P(K_{0}=k)$ and $p_{k}(t):=P(K_{t}=k|K_{0}=k)$.
 Proposition 13.
-For an IID process $(K_{t})_{t\in\mathbb{Z}}$, suppose that
-$\sum_{k}{\bf 1}{\left\{p_{k}\geq p\right\}}\leq C_{1}p^{-\beta}$ (65)
-for some $\beta\in(0,1)$, $C_{1}>0$, and all $p>0$. Then
-$\Delta V(t)\leq C_{2}t^{\beta-1},$ (66)
-where $C_{2}:=\Gamma(1-\beta)C_{1}$.
+For an IID process $(K_{t})_{t\in\mathbb{Z}}$ over natural numbers, suppose that
+$\sum_{k\in\mathbb{N}}{\bf 1}{\left\{p_{k}\geq p\right\}}\leq C_{0}p^{-\beta}$ (66)
+for some $\beta\in(0,1)$, $C_{0}>0$, and all $p>0$. Then
+$\Delta V(t)\leq C_{1}t^{\beta-1},$ (67)
+where $C_{1}:=\Gamma(1-\beta)C_{0}$.
 Proof.
-We may bound
-$\Delta V(t)$ $=\frac{V(t+1|1)}{t+1}=\sum_{k}p_{k}(1-p_{k})^{t}$
-$\leq\sum_{k}\int_{0}^{p_{k}}(1-p)^{t}dp=\int_{0}^{1}\left(\sum_{k}{\bf 1}{\left\{p_{k}\geq p\right\}}\right)(1-p)^{t}dp$
-$\leq C_{1}\int_{0}^{1}p^{-\beta}(1-p)^{t}dp.$ (67)
+By (65) and (62), we may bound
+$\Delta V(t)$ $=\frac{V(t+1|1)}{t+1}=\sum_{k\in\mathbb{N}}p_{k}(1-p_{k})^{t}$
+$\leq\sum_{k\in\mathbb{N}}\int_{0}^{p_{k}}(1-p)^{t}dp=\int_{0}^{1}\left(\sum_{k\in\mathbb{N}}{\bf 1}{\left\{p_{k}\geq p\right\}}\right)(1-p)^{t}dp$
+$\leq C_{0}\int_{0}^{1}p^{-\beta}(1-p)^{t}dp.$ (68)
 Further evaluation yields
 $\int_{0}^{1}(1-p)^{t}p^{-\beta}dp$ $=\frac{\Gamma(t+1)\Gamma(1-\beta)}{\Gamma(t+2-\beta)}\leq\Gamma(1-\beta)t^{\beta-1}.$ (69)
-The reverse bound can be demonstrated for a larger class of processes. For this aim, let us define a suitable mixing coefficient.
-Definition 5.
-For a stationary process $(K_{t})_{t\in\mathbb{Z}}$, denote the mixing coefficient
-$\Psi:=\sup_{t,k}\frac{\sum_{i=1}^{t}\sum_{j=1}^{t}P(K_{i}=k,K_{j}=k)-tp_{k}}{t(t-1)p_{k}^{2}}.$ (69)
-In particular, for an IID process, we have $\Psi=1$.
-For a stationary ergodic process, we have $\lim_{t\to\infty}t^{-1}\sum_{i=1}^{t}P(K_{0}=k,K_{i}=k)=p_{k}^{2}$ [26, Theorem 4.9].
-For a comprehensive survey of various mixing coefficients and mixing conditions, we refer to [12].
-For our purposes, it is sufficient to assume that $\Psi$ is finite.
+The reverse bound can be demonstrated for a larger class of processes.
 Proposition 14.
-For a stationary process $(K_{t})_{t\in\mathbb{Z}}$ with $\Psi<\infty$, suppose that
-$\sum_{k}p_{k}{\bf 1}{\left\{p_{k}\leq p\right\}}\geq C_{3}p^{1-\beta}$ (70)
-holds for some $\beta\in(0,1)$, $C_{3}>0$, and all $p>0$. Then we have
+For a stationary process $(K_{t})_{t\in\mathbb{Z}}$ over natural numbers, suppose that
+$\sum_{k\in\mathbb{N}}p_{k}{\bf 1}{\left\{p_{k}\leq p\right\}}$ $\geq C_{2}p^{1-\beta},$ (70)
+$\frac{p_{k}(t)}{p_{k}}$ $\leq C_{3}.$ (71)
+holds for some $\beta\in(0,1)$, $C_{2},C_{3}>0$, all $k,t\in\mathbb{N}$, and all $p>0$. Then we have
 $\Delta V(t)\geq C_{4}t^{\beta-1},$ (72)
-where $C_{4}:=(4\Psi)^{\beta-1}C_{3}/2$.
+where $C_{4}:=(4C_{3})^{\beta-1}C_{2}/2$.
 Proof.
@@
 $\Delta V(t)$ $\geq\frac{V(2t|1)}{2t}=\frac{1}{2t}\sum_{i=1}^{2t}P\left(K_{i}\not\in\mathcal{K}_{1}^{i-1}\cup\mathcal{K}_{i+1}^{2t}\right)$
-$=\sum_{k}\frac{1}{2t}\sum_{i=1}^{2t}P\left(K_{i}=k\not\in\mathcal{K}_{1}^{i-1}\cup\mathcal{K}_{i+1}^{2t}\right)$
-$=\sum_{k}\max\left\{0,\frac{1}{2t}\sum_{i=1}^{2t}P\left(K_{i}=k\not\in\mathcal{K}_{1}^{i-1}\cup\mathcal{K}_{i+1}^{2t}\right)\right\}$
-$\geq\sum_{k}\max\left\{0,2p_{k}-\frac{1}{2t}\sum_{i=1}^{2t}\sum_{j=1}^{2t}P(K_{i}=k,K_{j}=k)\right\}$
-$\geq\sum_{k}p_{k}\max\left\{0,1-\Psi(2t-1)p_{k}\right\}\geq\sum_{k}p_{k}\max\left\{0,1-2\Psi tp_{k}\right\}$
-$\geq\frac{1}{2}\sum_{k}p_{k}{\bf 1}{\left\{p_{k}\leq\frac{1}{4\Psi t}\right\}}\geq\frac{C_{3}(4\Psi t)^{\beta-1}}{2}.$ (72)
-Now we can show that conditions (65) and (70) hold under a broadly understood Zipf law.
+$=\sum_{k\in\mathbb{N}}\frac{1}{2t}\sum_{i=1}^{2t}P\left(K_{i}=k\not\in\mathcal{K}_{1}^{i-1}\cup\mathcal{K}_{i+1}^{2t}\right)$
+$=\sum_{k\in\mathbb{N}}\max\left\{0,\frac{1}{2t}\sum_{i=1}^{2t}P\left(K_{i}=k\not\in\mathcal{K}_{1}^{i-1}\cup\mathcal{K}_{i+1}^{2t}\right)\right\}$
+$\geq\sum_{k\in\mathbb{N}}\max\left\{0,2p_{k}-\frac{1}{2t}\sum_{i=1}^{2t}\sum_{j=1}^{2t}P(K_{i}=k,K_{j}=k)\right\}$
+$\geq\sum_{k\in\mathbb{N}}p_{k}\max\left\{0,1-C_{3}(2t-1)p_{k}\right\}\geq\sum_{k\in\mathbb{N}}p_{k}\max\left\{0,1-2C_{3}tp_{k}\right\}$
+$\geq\frac{1}{2}\sum_{k\in\mathbb{N}}p_{k}{\bf 1}{\left\{p_{k}\leq\frac{1}{4C_{3}t}\right\}}\geq\frac{C_{2}(4C_{3}t)^{\beta-1}}{2}.$ (73)
+We can show that conditions (66) and (70) hold under a broadly understood Zipf law.
 Example 1.
 Consider an approximate Zipf law
-$C_{5}k^{-\alpha}\leq p_{k}\leq C_{6}k^{-\alpha}$ (73)
-for some $\alpha>1$, $C_{5},C_{6}>0$, and all $k\in\mathbb{N}$. Then for $\beta=1/\alpha$, we have
-$\sum_{k}{\bf 1}{\left\{p_{k}\geq p\right\}}$ $\leq\int_{0}^{\infty}{\bf 1}{\left\{C_{6}k^{-\alpha}\geq p\right\}}dk$
+$C_{5}k^{-1/\beta}\leq p_{k}\leq C_{6}k^{-1/\beta}$ (74)
+for some $\beta\in(0,1)$, $C_{5},C_{6}>0$, and all $k\in\mathbb{N}$. Then we have
+$\sum_{k\in\mathbb{N}}{\bf 1}{\left\{p_{k}\geq p\right\}}$ $\leq\int_{0}^{\infty}{\bf 1}{\left\{C_{6}k^{-1/\beta}\geq p\right\}}dk$
 $=\int_{0}^{\infty}{\bf 1}{\left\{k\leq\left(\frac{p}{C_{6}}\right)^{-\beta}\right\}}dk=\left(\frac{p}{C_{6}}\right)^{-\beta},$ (75)
-$\sum_{k}p_{k}{\bf 1}{\left\{p_{k}\leq p\right\}}$ $\geq\int_{1}^{\infty}C_{5}k^{-\alpha}{\bf 1}{\left\{C_{6}k^{-\alpha}\leq p\right\}}dk$
-$=\int_{1}^{\infty}C_{5}k^{-\alpha}{\bf 1}{\left\{k\geq\left(\frac{p}{C_{6}}\right)^{-\beta}\right\}}dk=\frac{C_{5}}{1/\beta-1}\left(\frac{p}{C_{6}}\right)^{1-\beta}.$ (75)
-Thus conditions (65) and (70) hold for some constants $C_{1},C_{3}>0$.
-In consequence, an approximate Zipf law (73) implies the differential Heaps law (66) if the process is IID and it implies the differential Heaps law (71) if the process is stationary and $\Psi<\infty$.
+$\sum_{k\in\mathbb{N}}p_{k}{\bf 1}{\left\{p_{k}\leq p\right\}}$ $\geq\int_{1}^{\infty}C_{5}k^{-1/\beta}{\bf 1}{\left\{C_{6}k^{-1/\beta}\leq p\right\}}dk$
+$=\int_{1}^{\infty}C_{5}k^{-1/\beta}{\bf 1}{\left\{k\geq\left(\frac{p}{C_{6}}\right)^{-\beta}\right\}}dk=\frac{C_{5}}{1/\beta-1}\left(\frac{p}{C_{6}}\right)^{1-\beta}.$ (76)
+Thus conditions (66) and (70) hold for some constants $C_{0},C_{2}>0$.
+In consequence, an approximate Zipf law (74) implies the differential Heaps law (67) if the process is IID and, combined with (71), it implies the differential Heaps law (72) if the process is stationary.
+We suppose that the above results can be somewhat generalized by applying or developing classical results by Karlin [54].
 3.2 Heaps‘ law implies Hilberg‘s law
 Now we will formalize the remark made in Section 2.4.
-Namely, we will use the correspondence that the expected cardinality applies to arbitrary (stationary) processes as the Shannon entropy applies to (stationary) Santa Fe processes.
+Namely, we will use the correspondence that the expected cardinality applies to arbitrary (stationary) processes analogously as the Shannon entropy applies to (stationary) Santa Fe processes.
 In particular, the differential Heaps law (72) for a stationary narration implies a differential Hilberg law.
@@
 $\lim_{k\to\infty}H(K_{t+1}|K_{-k}^{t})=H(K_{t+1}|K_{-\infty}^{t})$ (83)
-holds in particular if the alphabet of process $(K_{t})_{t\in\mathbb{Z}}$ is finite (see [26, Sections 5.3 and 5.4] on a generalization of Shannon information measures to arbitrary $\sigma$-fields) or if $H(K_{t})<\infty$ (this follows by approximating $K_{t}$ by finite-alphabet random variables).
+holds in particular if the alphabet of process $(K_{t})_{t\in\mathbb{Z}}$ is finite (see [28, Sections 5.3 and 5.4] on a generalization of Shannon information measures to arbitrary $\sigma$-fields) or if $H(K_{t})<\infty$ (this follows by approximating $K_{t}$ by finite-alphabet random variables).
 Now suppose that process $(K_{t})_{t\in\mathbb{Z}}$ is stationary and $H(K_{t})<\infty$.
@@
 4 Conclusion
-Formalizing prior results in machine learning [50, 61, 64, 14, 70, 71] and extending earlier results in information theory and quantitative linguistics [23, 25, 26, 28], we have derived of a chain of deductive implications that connect Zipf‘s law to neural scaling.
-By isolating the discrete steps that go through Heaps‘ law and Hilberg‘s hypothesis and by giving explicit assumptions needed for each step, we have attempted to clarify which aspects of natural language properties are responsible for the power-law behavior observed in foundation models.
-Our results show that once vocabulary growth exhibits a power-law growth and once block entropy inherits this scaling, then the constraints imposed by limited data, parameters, and compute naturally produce a generic neural scaling law.
-Our theoretical derivation of underparameterization as the optimal regime contrasts sharply with the empirical success of overparameterized models [51].
+Formalizing prior results in machine learning [52, 65, 68, 15, 74, 75] and extending earlier results in information theory and quantitative linguistics [25, 27, 28, 30], we have derived of a deductive chain that connects Zipf‘s law to neural scaling.
+By isolating the discrete steps that go through Heaps‘ law and Hilberg‘s hypothesis and by giving explicit assumptions needed for each step, we have attempted to clarify which aspects of natural language are responsible for the power-law behavior observed in foundation models.
+Our results show that once vocabulary growth exhibits a power-law growth and once block entropy inherits this scaling, then the constraints imposed by limited data, parameters, and compute produce the neural scaling law.
+Our theoretical derivation of underparameterization as the optimal regime contrasts sharply with the empirical success of overparameterized models [53, 50, 76].
 This discrepancy suggests that naive parameter counts may overestimate the effective capacity of real-valued neural networks, as larger models may use fewer digits of the binary expansion per real-valued parameter.
@@
 Likewise, our purely information-theoretic modeling of compute—via entropy constraints—offers only a coarse abstraction of training dynamics.
-We suppose that a refined treatment may require tools from resource-bounded algorithmic information theory.
+We suppose that a refined treatment may rest on resource-bounded algorithmic information theory [59, 60].
 Finally, the reduction of the neural scaling law to Hilberg‘s hypothesis shifts attention to the deeper question of why power-law scaling of entropy arises in natural language at all.
-One simple possibility is that Hilberg‘s law emerges for Zipfian vocabularies with word shapes that are sufficiently arbitrary or algorithmically random, as suggested by the Santa Fe decomposition (7).
+One simple possibility is that Hilberg‘s law is equvalent to Zipf‘s law if word shapes are sufficiently arbitrary or algorithmically random, as suggested by the Santa Fe decomposition (7).
 Establishing such a connection rigorously requires a more detailed analysis of universal coding, extending those of works [27, 28, 29].
 We hope that the framework developed here provides a baseline for such investigations and that future work will refine the idealizations that we have identified.
-AI assistance
-Several paragraphs in the introduction and in the conclusion have been co-drafted by the author and ChatGPT and critically post-edited by the author. No reference was suggested by the AI assistant.
-Funding
+Acknowledgments and Disclosure of Funding
+Several paragraphs in the Introduction and in the Conclusion were drafted in a dialog between the author and ChatGPT (https://chatgpt.com/) and critically post-edited by the author.
+Subsequently, the article was reviewed by the Stanford Agentic Reviewer (https://paperreview.ai/) and its suggestions were partially followed by the author.
 This work received no external funding.